### v1

# Global small data weak solutions of 2-D semilinear wave equations with scale-invariant damping, III

Abstract
For the $2$-D semilinear wave equation with scale-invariant damping $\square u+\frac{\mu}{t}\partial_{t}u=|u|^{p}$, where $t\geq 1$, $\mu>0$ and $p>1$, it is conjectured that the global small data weak solution $u$ exists when $p>p_{s}(2+\mu)=\frac{\mu+3+\sqrt{\mu^{2}+14\mu+17}}{2(\mu+1)}$ for $0<\mu\leq 2$ and $p>p_{f}(2)=2$ for $\mu\geq 2$. In our previous papers, the global small solution $u$ has been obtained for $p>p_{s}(2+\mu)$ and $20<\mu<$ 2 but $\mu\not=1$. In the present paper, by the vector field method together with the delicate analysis on the Bessel functions, we will show the global existence of small solution $u$ for $p>2$ and $\mu\geq 2$. In forthcoming paper, for $\mu=1$ and $p>p_{s}(2+\mu)=p_{s}(3)=1+\sqrt{2}$, the global solution $u$ is also obtained. Therefore, this open question has been solved completely.
Keywords. Scale-invariant damping, global existence, Bessel function, vector field,
Klainerman-Sobolev inequality
2010 Mathematical Subject Classification 35L70, 35L65, 35L67
1 Introduction
In this paper, we continue to study the 2-D semilinear wave equation with scale-invariant damping
[…]
where $\mu>0$, $p>1$, $t\geq 1$, $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, $\partial_{i}=\partial_{x_{i}}$ ($i=1,2$), $\square=\partial_{t}^{2}-\Delta$ with $\Delta=\partial_{1}^{2}+\partial_{2}^{2}$, $u_{k}\in C_{0}^{\infty}(\mathbb{R}^{2})$ and $\operatorname{supp}u_{k}\in B(0,1)$ ($k=0,1$), $\varepsilon>0$ is small. The equation in (1.1) is also called as the semilinear Euler-Poisson-Darboux equation (see [1] and the references therein). For more detailed physical backgrounds on the linear wave operator $\square+\frac{\mu}{t}\partial_{t}$, one can be referred to [19] and so on.
Indicate the Strauss index $p_{s}(z)=\frac{z+1+\sqrt{z^{2}+10z-7}}{2(z-1)}$ ($z>1$) as a positive root of the quadratic algebraic equation $(z-1)p^{2}-(z+1)p-2=0$ (see [28] for the Strauss index $p_{s}(n)=\frac{n+1+\sqrt{n^{2}+10n-7}}{2(n-1)}$ of the semilinear wave equation $\square u=|u|^{p}$), and $p_{f}(n)=1+\frac{2}{n}$ as the Fujita index (see [9] for the Fujita index $p_{f}(n)=1+\frac{2}{n}$ of the semilinear parabolic equation $\partial_{t}u-\Delta u=|u|^{p}$). In terms of [3] and [15], there is an interesting open question on (1.1) as follows:
Open question (A). For 2-D problem (1.1),
(A1) when $20<\mu<$ 2 and $p>p_{s}(2+\mu)$, the small solution $u$ exists globally;
(A2) when $\mu\geq 2$ and $p>p_{f}(2)=2$, there exists a global solution $u$.
It is pointed out that the blowup results of problem (1.1) have been established for $1<p\leq\max\{p_{s}(2+\mu),p_{f}(2)\}$ (see [14], [18], [25]-[27], [29] and [30]). On the other hand, in our previous paper [19] and [11], (A1) has been proved except $\mu=1$. Note that (A2) with $\mu\geq 3$ or $\mu=2$ holds (see [2, 3]). Therefore, in order to solve (A2), we only require to study the case of $32<\mu<$ 3 in (1.1). Our main result is stated as follows.
Theorem 1.1.
For $32<\mu<$ 3 and $p>p_{f}(2)=2$, there exists a small constant $\varepsilon_{0}>0$ such that when $0<\varepsilon<\varepsilon_{0}$, (1.1) admits a unique global solution $u\in C\left([1,\infty);H^{2}\right)\cap C^{1}\left([1,\infty);H^{1}\right)\cap C^{2}\left([1,\infty);L^{2}\right)$.
Remark 1.1.
In our forthcoming paper [12], (A1) with $\mu=1$ and $p>p_{s}(2+\mu)=1+\sqrt{2}$ will be shown by applying the Bessel function tools and deriving some suitable time-decay or spacetime-decay estimates of weak solution $u$ as in [1], [11] and [19]. Collecting the results in [19] and [11] (for $20<\mu<$ 2 but $\mu\not=1$, $p>p_{s}(2+\mu)$), [12] (for $\mu=1$, $p>p_{s}(2+\mu)=1+\sqrt{2}$) together with Theorem 1.1 in the paper (for $32<\mu<$ 3 and $p\geq 2$) , [3, Theorem 2] (for $\mu=2$ and $p\geq 2$) and [2, Theorem 2] (for $\mu\geq 3$ and $p\geq 2$), then the Open question (A) has been solved completely.
Remark 1.2.
For the $n$-dimensional ($n\geq 3$) semilinear wave equation with scale-invariant damping
[…]
where $\square=\partial_{t}^{2}-(\partial_{1}^{2}+\cdot\cdot\cdot+\partial_{n}^{2})$, $\mu\in(0,1)\cup(1,2)$ and $p>p_{s}(n+\mu)=\frac{n+\mu+1+\sqrt{(n+\mu)^{2}+10(n+\mu)-7}}{2(n+\mu-1)}$, as pointed in Remark 1.12 of [11], the global existence of $u$ can be analogously shown by the methods in [19] and [11]. In addition, for $\mu=1$, $n\geq 4$ and $p>p_{s}(n+1)=\frac{n+2+\sqrt{n^{2}+12n+4}}{2n}$, the authors in [10] establish the global solution $u$ to problem (1.2).
Remark 1.3.
We emphasize that the series of methods developed in [19], [11] and the present paper will be very helpful in showing the global solvability of (1.2) when $\mu>0$ and $p>\max\{p_{s}(n+\mu),p_{f}(n)\}$.
In order to prove Theorem 1.1, let us recall the important vector field method in studying the global small data solution problem of $n-$dimensional ($n\geq 2$) quasilinear wave equation $\square\phi+\sum_{i,j,k=0}^{n}g_{ij}^{k}\partial_{ij}^{2}\phi\partial_{k}\phi=0$ (when $n=2,3$, the related null condition holds, namely, $\sum_{i,j,k=0}^{n}g_{ij}^{k}\xi_{i}\xi_{j}\xi_{k}$ $\equiv 0$ for $\xi_{0}=-1$ and $(\xi_{1},\cdot\cdot\cdot,\xi_{n})\in S^{n}$). Define the vector field $Z\in\{\partial,L_{0},L_{j},\Omega_{ij}:1\leq i<j\leq n\}$, where $\partial=(\partial_{0},\partial_{1},\ldots,\partial_{n})$ $=(\partial_{t},\partial_{x_{1}},\ldots,\partial_{x_{n}})$, $L_{0}=t\partial_{t}+x_{1}\partial_{1}+\cdots+x_{n}\partial_{n}$, $L_{j}=t\partial_{j}+x_{j}\partial_{t}$ and $\Omega_{ij}=x_{i}\partial_{j}-x_{j}\partial_{i}$. Then one easily has the following commutator properties
$[\square,\partial]=0,\quad[\square,L_{j}]=0,\quad[\square,\Omega_{ij}]=0,\quad[\square,L_{0}]=2\square.$
Based on this, it holds that for $Z^{I}\phi$ ($|I|\leq N_{0}$ with $N_{0}\in\mathbb{N}$ being suitably large),
$\square Z^{I}\phi+\sum_{i,j,k=0}^{n}\sum_{|I^{\prime}|+|I^{\prime\prime}|\leq|I|}C_{ij}^{k}\partial_{ij}^{2}Z^{I^{\prime}}\phi\partial_{k}Z^{I^{\prime\prime}}\phi=0.$ (1.3)
Therefore, it follows from (1.3) that $\{Z^{I}\phi\}_{|I|\leq N_{0}}$ consist of a closed nonlinear wave system and then $\|Z^{I}\partial\phi\|_{L_{x}^{2}(\mathbb{R}^{n})}$ can be obtained by the standard energy estimate. By the crucial Klainerman-Sobolev inequality
$|\partial\phi(t,x)|\leq\frac{C}{(1+|t-r|)^{\frac{1}{2}}(1+t)^{\frac{n-1}{2}}}\displaystyle\sum_{|J|\leq[\frac{n}{2}]+1}\|Z^{J}\partial\phi(t,x)\|_{L_{x}^{2}(\mathbb{R}^{n})},$ (1.4)
where $r=|x|=\sqrt{x_{1}^{2}+\cdot\cdot\cdot+x_{n}^{2}}$, then the time-decay and spacetime-decay rates of $\partial\phi$ are obtained. Making use of (1.4) and delicate weighted energies, a lot of interesting results on the global existence of the small data solution $\phi$ have been established. One can be referred to [4]-[6], [7]-[8], [13], [16]-[17], [20]-[21] and [22]-[23].
Note that for 2-D equation $\square u+\frac{2}{t}\,\partial_{t}u=|u|^{p}$, by the transformation $u=tv$, then the resulting semilinear wave equation $\square v=t^{-(p-1)}|v|^{p}$ can be obtained. By the standard vector field method and the energy estimate as in [21], the authors in Theorem 2 of [3] show that the small data solution $u\in C([1,\infty),H^{2}(\mathbb{R}^{2}))\cap C^{1}([1,\infty),H^{1}(\mathbb{R}^{2}))\cap C^{2}([1,\infty),L^{2}(\mathbb{R}^{2}))$ exists globally. However, for the general 2-D equation $\square u+\frac{\mu}{t}\,\partial_{t}u=|u|^{p}$ with $32<\mu<$ 3, it is difficult to find a suitable transformation of the unknown function $u$ such that a related semilinear wave equation is derived. Moreover, the commutator $[\square+\frac{\mu}{t}\,\partial_{t},Z]$ can not be expressed an efficient linear combination of $\square+\frac{\mu}{t}\,\partial_{t}$ and $Z$. Indeed, letting $w_{1}=\partial_{t}u$ and $w_{2}=L_{1}u$, then one has from $\square u+\frac{\mu}{t}\,\partial_{t}u=|u|^{p}$ that
$\square w_{1}+\frac{\mu}{t}\,\partial_{t}w_{1}-\frac{\mu}{t^{2}}w_{1}=\partial_{t}(|u|^{p})$ (1.5)
and
$\square w_{2}+\frac{\mu}{t}\,\partial_{t}w_{2}-\frac{\mu x_{1}}{t^{2}}w_{1}=L_{1}(|u|^{p}).$ (1.6)
This means that the linear parts in the nonlinear equations (1.5) and (1.6) become more complicated Euler-Poisson-Darboux-Klein-Gordon’s type models. To overcome this essential difficulty, we will take the following measures.
$\bullet$ By the Bessel function and Fourier analysis tools, through the explicit expression of solution to the 2-D linear equation
$\square v+\frac{\mu}{t}\,\partial_{t}v=F(t,x),\quad(v,\partial_{t}v)(1,x)=(v_{0},v_{1})(x),$ (1.7)
the required estimates of $v$ in some suitable norms are established. We will utilize the following norms for $v$:
$\|v\|_{Z,s,(p,q)}=\sum_{|\alpha|\leq s}\left\|Z^{\alpha}v\right\|_{(p,q)},\quad\|v\|_{Z,s,p}=\|v\|_{Z,s,(p,p)},\quad\|v\|_{Z,s,\infty}=\sum_{|\alpha|\leq s}\left\|Z^{\alpha}v\right\|_{\infty},$ (1.8)
where $s\in\mathbb{N}_{0}$, and $\|v\|_{(p,q)}:=\|v(r\omega)r^{\frac{1}{p}}\|_{L^{p}([0,+\infty);L^{q}(S^{1}))}$ for $p,q\in[1,\infty]$ and $\omega=\frac{x}{r}\in S^{1}$. It is pointed out that the norms in (1.8) and in higher space dimensions are firstly introduced in [21] for studying the lower bounds of the lifespan of classical solutions to the Cauchy problem for fully nonlinear wave equation $\square u=F(u,\partial u,\partial^{2}u)$, where the space dimensions $n\geq 3$ and the smooth nonlinearity $F(u,\partial u,\partial^{2}u)=O(|(u,\partial u,\partial^{2}u)|^{1+\lambda})$ ($\lambda\in\mathbb{N}$).
$\bullet$ Introducing the function space $X(T)=\{u(t,x)\in X(T):\sup_{t\in[1,T]}(t^{-(\delta-1)}\|u\|_{Z,1,2}+t\|\partial u\|_{Z,1,2})<\infty\}$ and the norm $\|u\|_{X(T)}:=\sup_{t\in[1,T]}(t^{-(\delta-1)}\|u\|_{Z,1,2}+t\|\partial u\|_{Z,1,2})$, where $T>1$ is any fixed number, $\delta>0$ is a suitable constant. On the other hand, it follows from (1.1) that one may define such a nonlinear mapping
$\mathcal{N}u(t,x)=\varepsilon\Psi_{0}(t,1,D)u_{0}(x)+\varepsilon\Psi_{1}(t,1,D)u_{1}(x)+\int_{1}^{t}\Psi_{1}(t,\tau,D)|u(\tau,x)|^{p}\mathrm{d}\tau,$ (1.9)
where the symbols of the pseudodifferential operators $\Psi_{0}(t,1,D)$ and $\Psi_{1}(t,1,D)$ are given in (2.4) and (2.5) of Section 2 below. Based on the key estimates for (1.8), we can show that the mapping $\mathcal{N}$ in (1.9) is contractible in a closed subspace of $X(T)$. Therefore, by the fixed point principle, problem (1.1) admits a global small solution $u\in C\left([1,\infty);H^{2}\right)\cap C^{1}\left([1,\infty);H^{1}\right)\cap C^{2}\left([1,\infty);L^{2}\right)$.
This paper is organized as follows. In Section 2, by the first kind of Bessel function and Hankel function, the explicit expression of the solution $v$ is given for the linear homogeneous equation $\square v+\frac{\mu}{t}\,\partial_{t}v=0$ with $(v,\partial_{t}v)(\tau,x)=(v_{0}(x),v_{1}(x))$. In Section 3, a series of time-decay estimates for the solution $v$ of $\square v+\frac{\mu}{t}\,\partial_{t}v=0$ are derived under the action of the vector field $Z\in\{\partial,L_{0},L_{j},\Omega_{12}:1\leq j\leq 2\}$. In addition, in Section 4, the time-decay estimates for the linear inhomogeneous equation $\square w+\frac{\mu}{t}\,\partial_{t}w=F$ with $(w,\partial_{t}w)(1,x)=(0,0)$ are established. In Section 5, based on the estimates in Sections 3-4, we complete the proof of Theorem 1.1 by using Duhamel’s principle and contraction mapping principle.
Below, we will use the following notations:
$\bullet$ For the nonnegative quantities $f$ and $g$, $f\lesssim g$ means $f\leq Cg$, where the generic constant $C>0$ is independent of $\varepsilon$;
$\bullet$ For the nonnegative quantities $f$ and $g$, $f\sim g$ means $g\lesssim f\lesssim g$;
$\bullet$ $\partial=(\partial_{t},\partial_{1},\partial_{2})$, $L_{0}=t\partial_{t}+x_{1}\partial_{1}+x_{2}\partial_{2}$, $L_{j}=t\partial_{j}+x_{j}\partial_{t}$ ($j=1,2$) and $\Omega_{12}=x_{1}\partial_{2}-x_{2}\partial_{1}$;
$\bullet$ $Z\in\{\partial,L_{0},L_{j},\Omega_{12}:1\leq j\leq 2\}$;
$\bullet$ For $p,q\in[1,\infty]$, $\|v\|_{(p,q)}:=\|v(r\omega)r^{\frac{1}{p}}\|_{L^{p}([0,+\infty);L^{q}(S^{1}))}$ with $\omega=\frac{x}{r}\in S^{1}$;
$\bullet$ For $s\in\mathbb{N}_{0}$ and $p,q\in[1,\infty]$,
$\|v\|_{Z,s,(p,q)}=\sum_{|\alpha|\leq s}\left\|Z^{\alpha}v\right\|_{(p,q)},\quad\|v\|_{Z,s,p}=\|v\|_{Z,s,(p,p)},\quad\|v\|_{Z,s,\infty}=\sum_{|\alpha|\leq s}\left\|Z^{\alpha}v\right\|_{\infty},$
where $\|v\|_{(p,q)}:=\|v(r\omega)r^{\frac{1}{p}}\|_{L^{p}([0,+\infty);L^{q}(S^{1}))}$ and $\omega=\frac{x}{r}\in S^{1}$;
$\bullet$ For $\xi\in\mathbb{R}^{2}$ and $t\geq 1$, define
$A_{1}=\{\xi:|\xi|\geq 1\},\quad A_{2}=\{\xi:|\xi|\leq 1\leq t|\xi|\},\quad A_{3}=\{\xi:t|\xi|\leq 1\}.$
2 The expression of solution to 2-D equation $\square v+\frac{\mu}{t}\,\partial_{t}v=0$
In this section, we derive the expression of the solution $v$ to 2-D problem
[…]
where $32<\mu<$ 3 and $v_{i}(x)\in C_{0}^{\infty}(\mathbb{R}^{2})$ $(i=0,1)$.
Taking the Fourier transformation of $v(t,x)$ with respect to the variable $x$, then one has from (2.1) that
[…]
It follows from [32, Theorem 2.1] that the solution $\hat{v}(t,\xi)$ of (2.2) can be expressed as
$\hat{v}(t,\xi)=\Psi_{0}(t,\tau,\xi)\hat{v}_{0}(\xi)+\Psi_{1}(t,\tau,\xi)\hat{v}_{1}(\xi),$ (2.3)
where the multipliers $\Psi_{j}(t,\tau,\xi)(j=0,1)$ and its time derivatives are given by
$\Psi_{0}(t,\tau,\xi)=\frac{i\pi}{4}|\xi|\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho}^{+}(t|\xi|)&H_{\rho-1}^{+}(\tau|\xi|)\\ H_{\rho}^{-}(t|\xi|)&H_{\rho-1}^{-}(\tau|\xi|)\end{array}\right|,$ (2.4)
$\Psi_{1}(t,\tau,\xi)=-\frac{i\pi}{4}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho}^{+}(t|\xi|)&H_{\rho}^{+}(\tau|\xi|)\\ H_{\rho}^{-}(t|\xi|)&H_{\rho}^{-}(\tau|\xi|)\end{array}\right|,$ (2.5)
$\partial_{t}^{k}\Psi_{j}(t,\tau,\xi)=(-1)^{j}\frac{i\pi}{4}|\xi|^{1+k-j}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho-k}^{+}(t|\xi|)&H_{\rho-1+j}^{+}(\tau|\xi|)\\ H_{\rho-k}^{-}(t|\xi|)&H_{\rho-1+j}^{-}(\tau|\xi|)\end{array}\right|,\quad k,j=0,1,$ (2.6)
where $\rho=-\frac{\mu-1}{2}\in(-1,-\frac{1}{2})$, $H_{\nu}^{\pm}$ are the Hankel functions of order $\nu$ ($\nu\not\in\mathbb{Z}$), which can be expressed as (see [24, Section 10])
$H_{\nu}^{+}(z)=i\csc(\nu\pi)\left(e^{-\nu\pi i}J_{\nu}(z)-J_{-\nu}(z)\right),$ (2.7)
$H_{\nu}^{-}(z)=i\csc(\nu\pi)\left(J_{-\nu}(z)-e^{\nu\pi i}J_{\nu}(z)\right)$
with the first kind of Bessel function $J_{\gamma}(z)=(\frac{z}{2})^{\gamma}\sum_{k=0}^{\infty}(-1)^{k}\frac{z^{2k}}{4^{k}k!\Gamma(\gamma+k+1)}$ for $\gamma\in\mathbb{C}\setminus\mathbb{Z}$.
Therefore,
$\partial_{t}^{k}\Psi_{j}(t,\tau,\xi)=\frac{\pi}{2}\csc(\rho\pi)|\xi|^{1+k-j}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{cc}J_{-(\rho-1+j)}(\tau|\xi|)&J_{-(\rho-k)}(t|\xi|)\\ (-1)^{1+k-j}J_{\rho-1+j}(\tau|\xi|)&J_{\rho-k}(t|\xi|)\end{array}\right|,\quad k,j=0,1.$ (2.8)
Based on (2.8), we have
Lemma 2.1.
It holds that for $\rho=-\frac{\mu-1}{2}\in(-1,-\frac{1}{2})$,
$\partial_{t}^{2}\Psi_{0}(t,\tau,\xi)$ $=\frac{i\pi}{4}|\xi|^{3}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho-2}^{+}(t|\xi|)&H_{\rho-1}^{+}(\tau|\xi|)\\ H_{\rho-2}^{-}(t|\xi|)&H_{\rho-1}^{-}(\tau|\xi|)\end{array}\right|+t^{-1}\partial_{t}\Psi_{0}(t,\tau,\xi),$ (2.9)
$\partial_{t}^{2}\Psi_{1}(t,\tau,\xi)$ $=-\frac{i\pi}{4}|\xi|^{2}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho-2}^{+}(t|\xi|)&H_{\rho}^{+}(\tau|\xi|)\\ H_{\rho-2}^{-}(t|\xi|)&H_{\rho}^{-}(\tau|\xi|)\end{array}\right|+t^{-1}\partial_{t}\Psi_{1}(t,\tau,\xi).$ (2.10)
Meanwhile, one also has
$\partial_{t}^{2}\Psi_{0}(t,\tau,\xi)$ $=\frac{\pi}{2}|\xi|^{3}\frac{t^{\rho}}{\tau^{\rho-1}}\csc(\rho\pi)\left|\begin{array}[]{ll}J_{-(\rho-1)}(\tau|\xi|)&J_{-(\rho-2)}(t|\xi|)\\ -J_{\rho-1}(\tau|\xi|)&J_{\rho-2}(t|\xi|)\end{array}\right|+t^{-1}\partial_{t}\Psi_{0}(t,\tau,\xi),$ (2.11)
$\partial_{t}^{2}\Psi_{1}(t,\tau,\xi)$ $=\frac{\pi}{2}|\xi|^{2}\frac{t^{\rho}}{\tau^{\rho-1}}\csc(\rho\pi)\left|\begin{array}[]{ll}J_{-\rho}(\tau|\xi|)&J_{-(\rho-2)}(t|\xi|)\\ J_{\rho}(\tau|\xi|)&J_{\rho-2}(t|\xi|)\end{array}\right|+t^{-1}\partial_{t}\Psi_{1}(t,\tau,\xi).$ (2.12)
Proof.
Motivated by [32, Corollary 2.2], setting $r=t|\xi|$, then it follows from (2.2) and the assumption of $\hat{v}=r^{\rho}y(r)$ that $y(r)$ satisfies
$r^{2}y^{\prime\prime}(r)+ry^{\prime}(r)+\left(r^{2}-\rho^{2}\right)y(r)=0.$ (2.13)
Hence, by (2.13), (2.2) admits two independent solutions $\hat{v}_{+}(r)=r^{\rho}H_{\rho}^{+}(r)$ and $\hat{v}_{-}(r)=r^{\rho}H_{\rho}^{-}(r)$. Therefore,
$\Psi_{0}(t,\tau,\xi)=C_{01}(\tau,\xi)\hat{v}_{+}(t,\xi)+C_{02}(\tau,\xi)\hat{v}_{-}(t,\xi),$ (2.14)
$\Psi_{1}(t,\tau,\xi)=C_{11}(\tau,\xi)\hat{v}_{+}(t,\xi)+C_{12}(\tau,\xi)\hat{v}_{-}(t,\xi),$
where
$\hat{v}_{+}(t,\xi)=t^{\rho}|\xi|^{\rho}H_{\rho}^{+}(t|\xi|),\quad\hat{v}_{-}(t,\xi)=t^{\rho}|\xi|^{\rho}H_{\rho}^{-}(t|\xi|).$ (2.15)
In addition, it follows from (16)-(19) in [32] that
$\left(\begin{array}[]{cc}C_{01}&C_{11}\\ C_{02}&C_{12}\end{array}\right)=\frac{1}{\left|\begin{array}[]{cc}\hat{v}_{+}(\tau,\xi)&\hat{v}_{-}(\tau,\xi)\\ \hat{\partial}_{t}\hat{v}_{+}(\tau,\xi)&\partial_{t}\hat{v}_{-}(\tau,\xi)\end{array}\right|}\left(\begin{array}[]{cc}\partial_{t}\hat{v}_{-}(\tau,\xi)&-\hat{v}_{-}(\tau,\xi)\\ -\partial_{t}\hat{v}_{+}(\tau,\xi)&\hat{v}_{+}(\tau,\xi)\end{array}\right)$ (2.16)
$=$ $\frac{i\pi}{4\tau^{2\rho-1}|\xi|^{2\rho}}\left(\begin{array}[]{cc}\tau^{\rho}|\xi|^{\rho+1}H_{\rho-1}^{-}(\tau|\xi|)&-\tau^{\rho}|\xi|^{\rho}H_{\rho}^{-}(\tau|\xi|)\\ -\tau^{\rho}|\xi|^{\rho+1}H_{\rho-1}^{+}(\tau|\xi|)&\tau^{\rho}|\xi|^{\rho}H_{\rho}^{+}(\tau|\xi|)\end{array}\right).$
By $\nu H_{\nu}^{\pm}(z)+z(H_{\nu}^{\pm})^{\prime}(z)=zH_{\nu-1}^{\pm}(z)$ and (2.15), we arrive at
$\partial_{t}\hat{v}_{+}(t,\xi)=t^{\rho}|\xi|^{\rho+1}H_{\rho-1}^{+}(t|\xi|),~{}\partial_{t}\hat{v}_{-}(t,\xi)=t^{\rho}|\xi|^{\rho+1}H_{\rho-1}^{-}(t|\xi|),$ (2.17)
$\partial_{t}^{2}\hat{v}_{+}(t,\xi)=t^{\rho}|\xi|^{\rho+2}H_{\rho-2}^{+}(t|\xi|)+t^{\rho-1}|\xi|^{\rho+1}H_{\rho-1}^{+}(t|\xi|),$
$\partial_{t}^{2}\hat{v}_{-}(t,\xi)=t^{\rho}|\xi|^{\rho+2}H_{\rho-2}^{-}(t|\xi|)+t^{\rho-1}|\xi|^{\rho+1}H_{\rho-1}^{-}(t|\xi|).$
Substituting (2.16) and (2.17) into (2.14) yields
$\partial_{t}^{2}\Psi_{0}(t,\tau,\xi)$ $=\frac{i\pi}{4}|\xi|^{3}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho-2}^{+}(t|\xi|)&H_{\rho-1}^{+}(\tau|\xi|)\\ H_{\rho-2}^{-}(t|\xi|)&H_{\rho-1}^{-}(\tau|\xi|)\end{array}\right|+\frac{i\pi}{4}|\xi|^{2}\frac{t^{\rho-1}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho-1}^{+}(t|\xi|)&H_{\rho-1}^{+}(\tau|\xi|)\\ H_{\rho-1}^{-}(t|\xi|)&H_{\rho-1}^{-}(\tau|\xi|)\end{array}\right|$ (2.18)
$=\frac{i\pi}{4}|\xi|^{3}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho-2}^{+}(t|\xi|)&H_{\rho-1}^{+}(\tau|\xi|)\\ H_{\rho-2}^{-}(t|\xi|)&H_{\rho-1}^{-}(\tau|\xi|)\end{array}\right|+t^{-1}\partial_{t}\Psi_{0}(t,\tau,\xi)$
and
$\partial_{t}^{2}\Psi_{1}(t,\tau,\xi)$ $=-\frac{i\pi}{4}|\xi|^{2}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho-2}^{+}(t|\xi|)&H_{\rho}^{+}(\tau|\xi|)\\ H_{\rho-2}^{-}(t|\xi|)&H_{\rho}^{-}(\tau|\xi|)\end{array}\right|+t^{-1}\partial_{t}\Psi_{1}(t,\tau,\xi).$ (2.19)
Since $\rho$ is not an integer, one then has that from (2.7) and (2.18),
$\frac{i\pi}{4}|\xi|^{3}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho-2}^{+}(t|\xi|)&H_{\rho-1}^{+}(\tau|\xi|)\\ H_{\rho-2}^{-}(t|\xi|)&H_{\rho-1}^{-}(\tau|\xi|)\end{array}\right|$ (2.20)
$=\frac{i\pi}{4}|\xi|^{3}\frac{t^{\rho}}{\tau^{\rho-1}}\csc^{2}(\rho\pi)\left(e^{-(\rho-2)\pi i}J_{-(\rho-1)}(\tau|\xi|)J_{\rho-2}(t|\xi|)-J_{-(\rho-1)}(\tau|\xi|)J_{-(\rho-2)}(t|\xi|)\right.$
$\quad\left.-e^{\pi i}J_{\rho-1}(\tau|\xi|)J_{\rho-2}(t|\xi|)+e^{(\rho-1)\pi i}J_{\rho-1}(\tau|\xi|)J_{-(\rho-2)}(t|\xi|)\right)$
$\quad-\frac{i\pi}{4}|\xi|^{3}\frac{t^{\rho}}{\tau^{\rho-1}}\csc^{2}(\rho\pi)\left(e^{-(\rho-1)\pi i}J_{\rho-1}(\tau|\xi|)J_{-(\rho-2)}(t|\xi|)-e^{-\pi i}J_{\rho-1}(\tau|\xi|)J_{\rho-2}(t|\xi|)\right.$
$\quad\left.-J_{-(\rho-1)}(\tau|\xi|)J_{-(\rho-2)}(t|\xi|)+e^{(\rho-2)\pi i}J_{-(\rho-1)}(\tau|\xi|)J_{\rho-2}(t|\xi|)\right)$
$=\frac{i\pi}{4}|\xi|^{3}\frac{t^{\rho}}{\tau^{\rho-1}}\csc^{2}(\rho\pi)\left[-2i\sin(\rho\pi)\cdot\left(J_{\rho-1}(\tau|\xi|)J_{\rho-2}(t|\xi|)+J_{\rho-1}(\tau|\xi|)J_{-(\rho-2)}(t|\xi|)\right)\right]$
$=\frac{\pi}{2}|\xi|^{3}\frac{t^{\rho}}{\tau^{\rho-1}}\csc(\rho\pi)\left|\begin{array}[]{cc}J_{-(\rho-1)}(\tau|\xi|)&J_{-(\rho-2)}(t|\xi|)\\ -J_{\rho-1}(\tau|\xi|)&J_{\rho-2}(t|\xi|)\end{array}\right|.$
Similarly,
$-\frac{i\pi}{4}|\xi|^{2}\frac{t^{\rho}}{\tau^{\rho-1}}\left|\begin{array}[]{ll}H_{\rho-2}^{+}(t|\xi|)&H_{\rho}^{+}(\tau|\xi|)\\ H_{\rho-2}^{-}(t|\xi|)&H_{\rho}^{-}(\tau|\xi|)\end{array}\right|=\frac{\pi}{2}|\xi|^{2}\frac{t^{\rho}}{\tau^{\rho-1}}\csc(\rho\pi)\left|\begin{array}[]{ll}J_{-\rho}(\tau|\xi|)&J_{-(\rho-2)}(t|\xi|)\\ J_{\rho}(\tau|\xi|)&J_{\rho-2}(t|\xi|)\end{array}\right|.$ (2.21)
Therefore, (2.11) and (2.12) follow from (2.20) and (2.21), respectively. ∎
Let us recall some useful results on the asymptotic behaviors of $J_{\nu}(z)$ and $H_{\nu}^{\pm}(z)$ ($z>0$) for sufficiently large $z$ or small $z$. For details, one can be refered to [24, Section 10] or [31, §3.13, §3.52, §10.6 and §7.2].
Lemma 2.2.
It holds that
(i) for large $z\geq K>0$,
$|H_{\nu}^{\pm}(z)|\lesssim z^{-\frac{1}{2}};$ (2.22)
for small $z$ with $10<z\leq c<$ 1,
$\left|H_{\nu}^{\pm}(z)\right|\lesssim\begin{cases}z^{-|\nu|},&\nu\neq 0,\\ -\log z,&\nu=0.\end{cases}$ (2.23)
(ii) for large $z\geq K>0$,
$|J_{\nu}(z)|\lesssim z^{-\frac{1}{2}};$ (2.24)
for small $z$ with $10<z\leq c<$ 1,
$|J_{\nu}(z)|\lesssim z^{\nu}.$ (2.25)
3 Estimates of solution to 2-D homogeneous equation $\square v+\frac{\mu}{t}\,\partial_{t}v=0$
Let $v$ solve
[…]
where $32<\mu<$ 3 and $v_{i}(x)\in C_{0}^{\infty}(B(0,1))$ $(i=0,1)$.
Lemma 3.1.
For any $\varepsilon_{1}\in(0,1)$, there exists a constant $\delta(\varepsilon_{1})=\frac{2\varepsilon_{1}}{1+\varepsilon_{1}}>0$ such that
$\|v(t,\cdot)\|_{Z,1,2}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.2)
Proof.
By the definition of $\|v\|_{Z,s,(p,q)}$ with $s=1$ and $p=q=2$, we firstly establish
$\|v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{L^{2}(\mathbb{R}^{2})}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\|v_{0}+v_{1}\|_{(1+\varepsilon_{1},2)}.$ (3.3)
Note that $\hat{v}(t,\xi)=\Psi_{0}(t,1,\xi)\hat{v}_{0}(\xi)+\Psi_{1}(t,1,\xi)\hat{v}_{1}(\xi)$. In order to use the asymptotic behaviors of Bessel functions to estimate $v(t,x)$, we will decompose the frequency $\xi\in\mathbb{R}^{2}$ into the following three zones
$A_{1}=\{\xi:|\xi|\geq 1\},\quad A_{2}=\{\xi:|\xi|\leq 1\leq t|\xi|\},\quad A_{3}=\{\xi:t|\xi|\leq 1\}.$ (3.4)
Part 1. The analysis in zone $A_{1}$
By (2.6) with $\rho=-\frac{\mu-1}{2}$ and (2.22), one can obtain
$\left|\Psi_{0}(t,1,\xi)\right|\lesssim|\xi|t^{\rho}|\xi|^{-\frac{1}{2}}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}},$ (3.5)
$\left|\Psi_{1}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{-\frac{1}{2}}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|^{-1}.$
Then
$\left\|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}$ $\leq\left\|\Psi_{0}(t,1,\xi)\right\|_{L^{\infty}}\left\|\hat{v}_{0}\right\|_{L^{2}}+\left\||\Psi_{1}(t,1,\xi)|\hat{v}_{1}\right\|_{L^{2}}$ (3.6)
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\|(1+|\xi|^{2})^{-\frac{1}{2}}\hat{v}_{1}\|_{L^{2}}:=t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{L^{2}}+I.$
Next we treat the term $I$ in (3.6). By setting $x=ty$ and $\xi=\frac{\eta}{t}$, one has that for $32<\mu<$ 3 and $t\geq 1$,
$I$ $=t^{-\frac{\mu}{2}}\left(\int_{R^{2}}\frac{1}{1+|\xi|^{2}}\big{(}\int_{R^{2}}e^{-ix\cdot\xi}v_{1}(x)\mathrm{d}x\big{)}^{2}\mathrm{d}\xi\right)^{\frac{1}{2}}$ (3.7)
$=t^{-\frac{\mu}{2}+2}\left(\int_{R^{2}}\frac{1}{t^{2}+|\eta|^{2}}\big{(}\int_{R^{2}}e^{-iy\cdot\eta}v_{1}(ty)\mathrm{d}y\big{)}^{2}\mathrm{d}\eta\right)^{\frac{1}{2}}$
$\leq t^{-\frac{\mu}{2}+2}\left(\int_{R^{2}}\frac{1}{1+|\eta|^{2}}\big{(}\int_{R^{2}}e^{-iy\cdot\eta}v_{1}(ty)\mathrm{d}y\big{)}^{2}\mathrm{d}\eta\right)^{\frac{1}{2}}$
$\lesssim t\|v_{1}(ty)\|_{H^{-1}(\mathbb{R}^{2})}.$
Due to $\|V_{1}\|_{H^{-1}}=\sup_{v\in H^{1},v\neq 0}\frac{\left|\int_{\mathbb{R}^{2}}V_{1}(y)v(y)\mathrm{d}y\right|}{\|v\|_{H^{1}}}$ for $V_{1}(y)=v_{1}(ty)$, then it follows from Hölder’s inequality and Sobolev imbedding theorem that
$\|V_{1}v\|_{L^{1}}\lesssim\|V_{1}\|_{(q,2)}\|v\|_{\left(q^{\prime},2\right)}\lesssim\|V_{1}\|_{(q,2)}\|v\|_{q^{\prime}}\lesssim\|V_{1}\|_{(q,2)}\|v\|_{H^{1}}\lesssim t^{-\frac{2}{q}}\|v_{1}\|_{(q,2)}\|v\|_{H^{1}},$ (3.8)
where $q=1+\varepsilon_{1}$ with any fixed number $\varepsilon_{1}\in(0,1)$ and $\frac{1}{q}+\frac{1}{q^{\prime}}=1$. Hence,
$\|V_{1}\|_{H^{-1}}\lesssim t^{-\frac{2}{1+\varepsilon_{1}}}\|v_{1}\|_{(1+\varepsilon_{1},2)}.$
This, together with (3.7), yields
$I\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{1}\|_{(1+\varepsilon_{1},2)}$
and
$\left\|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{L^{2}}+t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{1}\|_{(1+\varepsilon_{1},2)}.$ (3.9)
Part 2. The analysis in zone $A_{2}$
For $\rho=-\frac{\mu-1}{2}\in(-1,-\frac{1}{2})$, it follows from (2.8), (2.24) and (2.25) that
$\left|\Psi_{0}(t,1,\xi)\right|\lesssim|\xi|t^{\rho}|\xi|^{\rho-1}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}},$ (3.10)
$\left|\Psi_{1}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{\rho}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}}.$
Due to $t|\xi|\geq 1$, then one can utilize $(t|\xi|)^{-\frac{\mu}{2}}\lesssim(1+t|\xi|)^{-1}$, $x=ty$ and $t\xi=\eta$ to obtain
$\left\|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}$ $\leq\left\|(1+t|\xi|)^{-1}(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}$ (3.11)
$=t\left(\int_{R^{2}}(\frac{1}{1+|\eta|})^{2}\big{(}\int_{R^{2}}e^{-iy\cdot\eta}(v_{0}+v_{1})(ty)\mathrm{d}y\big{)}^{2}\mathrm{d}\eta\right)^{\frac{1}{2}}$
$\lesssim t\|(v_{0}+v_{1})(ty)\|_{H^{-1}(\mathbb{R}^{2})}.$
Analogously to the treatment of (3.8), we then have
$\left\|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{0}+v_{1}\|_{(1+\varepsilon_{1},2)}.$ (3.12)
Part 3. The analysis in zone $A_{3}$
As in the zone $A_{2}$, one has
$\left|\Psi_{0}(t,1,\xi)\right|\lesssim|\xi|t^{\rho}|\xi|^{\rho-1}(t|\xi|)^{-\rho}=1,$ (3.13)
$\left|\Psi_{1}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{\rho}(t|\xi|)^{-\rho}=1.$
Then for $|\xi|\leq t|\xi|\leq 1$, as in (3.11), we arrive at
$\left\|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{3})}$ $\leq\left\|(1+t|\xi|)\times(1+t|\xi|)^{-1}(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}$ (3.14)
$\lesssim\left\|(1+t|\xi|)^{-1}(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}$
$\lesssim t\|(v_{0}+v_{1})(ty)\|_{H^{-1}(\mathbb{R}^{2})}$
$\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{0}+v_{1}\|_{(1+\varepsilon_{1},2)}.$
Collecting (3.9), (3.12), (3.14) and Parseval’s equality yields (3.3).
Next we prove (3.2). By the definition of $\|v\|_{Z,1,2}$, the proof procedure will be divided into the following four cases.
Case 1. The treatment of $\|\partial v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$
In this case, we shall establish the following estimate
$\|\partial v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.15)
To prove (3.15), we still decompose the space $\mathbb{R}_{\xi}^{2}$ into three zones as in (3.4).
Case 1.1. The analysis in zone $A_{1}$
It follows from (2.6) and (2.22) that
$\left|\partial_{t}\Psi_{0}(t,1,\xi)\right|\lesssim|\xi|^{2}t^{\rho}|\xi|^{-\frac{1}{2}}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|,$ (3.16)
$\left|\partial_{t}\Psi_{1}(t,1,\xi)\right|\lesssim|\xi|t^{\rho}|\xi|^{-\frac{1}{2}}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}.$
Then for $|\xi|\geq 1$ and $32<\mu<$ 3, as in (3.11), one has
$\left\|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}$ $\lesssim t^{-\frac{\mu}{2}}\left\||\xi|v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\||\xi|^{-1}|\xi|\hat{v}_{1}\|_{L^{2}}$ (3.17)
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}+1}\|(1+t|\xi|)^{-1}|\xi|\hat{v}_{1}\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}+2}\|\nabla v_{1}(ty)\|_{H^{-1}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{1}\|_{Z,1,(1+\varepsilon_{1},2)}.$
Analogously to the treatment of $\left\|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}$, we have that for $j=1,2$,
$\left\||\xi_{j}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{1}\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.18)
This, together with (3.17), yields
$\|\widehat{\partial v}\|_{L^{2}(A_{1})}\leq\left\|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}+\sum_{j=1}^{2}\left\||\xi_{j}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.19)
Case 1.2. The analysis in zone $A_{2}$
In this case, it follows from (2.8), (2.24) and (2.25) that
$\left|\partial_{t}\Psi_{0}(t,1,\xi)\right|\lesssim|\xi|^{2}t^{\rho}|\xi|^{\rho-1}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}}|\xi|\leq(t|\xi|)^{-\frac{\mu}{2}},$ (3.20)
$\left|\partial_{t}\Psi_{1}(t,1,\xi)\right|\lesssim|\xi|t^{\rho}|\xi|^{\rho}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}}|\xi|\leq(t|\xi|)^{-\frac{\mu}{2}}.$
Following (3.11), by the change of variables $\eta=t\xi$ and $x=ty$, we have
$\left\|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}$ $\leq\left\|(1+t|\xi|)^{-1}(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}$ (3.21)
$\lesssim t\|(v_{0}+v_{1})(ty)\|_{H^{-1}(\mathbb{R}^{2})}$
$\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{0}+v_{1}\|_{(1+\varepsilon_{1},2)}.$
Due to $|\xi_{j}|\leq|\xi|\leq 1(j=1,2)$, then
$\left\||\xi_{j}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{0}+v_{1}\|_{(1+\varepsilon_{1},2)}$ (3.22)
and
$\|\widehat{\partial v}\|_{L^{2}(A_{2})}\lesssim t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.23)
Case 1.3. The analysis in zone $A_{3}$
By $t|\xi|\leq 1$, one has from (2.8) and (2.25) that
$\left|\partial_{t}\Psi_{0}(t,1,\xi)\right|\lesssim|\xi|^{2}t^{\rho}[|\xi|^{-\rho+1}(t|\xi|)^{\rho-1}+|\xi|^{\rho-1}(t|\xi|^{1-\rho})]=|\xi|^{2}(t^{-\mu}+t)\leq t^{-1},$ (3.24)
$\left|\partial_{t}\Psi_{1}(t,1,\xi)\right|\lesssim|\xi|t^{\rho}[|\xi|^{-\rho}(t|\xi|)^{\rho-1}+|\xi|^{\rho}(t|\xi|)^{1-\rho}]=t^{-\mu}+t|\xi|^{2}\leq t^{-1}.$
Then
$\left\|\partial_{t}\hat{v}\right\|_{L^{2}(A_{3})}$ $\leq t^{-1}\left\|(1+t|\xi|)\times(1+t|\xi|)^{-1}(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}$ (3.25)
$\lesssim t^{-1}\left\|(1+t|\xi|)^{-1}(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}$
$\lesssim\|(v_{0}+v_{1})(ty)\|_{H^{-1}(\mathbb{R}^{2})}$
$\lesssim t^{-\frac{2}{1+\varepsilon_{1}}}\|v_{0}+v_{1}\|_{(1+\varepsilon_{1},2)}$
and
$\left\||\xi_{j}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{3})}\lesssim t^{-\frac{2}{1+\varepsilon_{1}}}\|v_{0}+v_{1}\|_{(1+\varepsilon_{1},2)}\quad\quad\text{for}~{}j=1,2.$ (3.26)
Combining (3.25) and (3.26) yields
$\|\widehat{\partial v}\|_{L^{2}(A_{3})}\lesssim t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.27)
Therefore, by Cases 1.1-1.3, we have
$\|\partial v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$ $=\|\widehat{\partial v}\|_{L^{2}(\mathbb{R}^{2})}\leq\|\widehat{\partial v}\|_{L^{2}(A_{1})}+\|\widehat{\partial v}\|_{L^{2}(A_{2})}+\|\widehat{\partial v}\|_{L^{2}(A_{3})}$ (3.28)
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$
Case 2. The treatment of $\|L_{0}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$
In this case, we start to show
$\|L_{0}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.29)
Note that
$\|L_{0}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$ $=\|t\partial_{t}\hat{v}+\widehat{x_{1}\partial_{1}v}+\widehat{x_{2}\partial_{2}v}\|_{L^{2}(\mathbb{R}^{2})}$ (3.30)
$\leq 2\|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}.$
Thus, in order to prove (3.29), it suffices to treat $\|t\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ and $\|\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ since the estimate of $\|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ has been derived in (3.3).
For $\|t\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$, it follows from (3.16) and (3.20) that for $32<\mu<$ 3,
$\left\|t\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}$ $\lesssim t^{-\frac{\mu}{2}}\left\|t|\xi|\hat{v}_{0}\right\|_{L^{2}}+t^{1-\frac{\mu}{2}}\||\xi|^{-1}|\xi|\hat{v}_{1}\|_{L^{2}}$ (3.31)
$\lesssim t^{-\frac{\mu}{2}}(\|L_{1}v_{0}\|_{L^{2}}+\|L_{2}v_{0}\|_{L^{2}})+\|(1+t|\xi|)^{-1}t|\xi|\hat{v}_{1}\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{1-\frac{2}{1+\varepsilon_{1}}}(\|t\partial_{1}v_{1}\|_{(1+\varepsilon_{1},2)}+\|t\partial_{2}v_{1}\|_{(1+\varepsilon_{1},2)})$
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}$
and
$\left\|t\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}\leq\left\|(1+t|\xi|)^{-1}t|\xi|(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}$ (3.32)
$\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}(\|t\partial_{1}(v_{0}+v_{1})+x_{1}\partial_{t}(v_{0}+v_{1})\|_{(1+\varepsilon_{1},2)}+\|t\partial_{2}(v_{0}+v_{1})+x_{2}\partial_{t}(v_{0}+v_{1})\|_{(1+\varepsilon_{1},2)})$
$\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{0}+v_{1}\|_{Z,1,(1+\varepsilon_{1},2)}.$
Note that from (3.25),
$\left\|t\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{3})}\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\|v_{0}+v_{1}\|_{(1+\varepsilon_{1},2)}.$ (3.33)
This, together with (3.31) and (3.32), derives
$\|t\partial_{t}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.34)
We next treat $\|\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ and further assert
$\|\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.35)
To prove (3.35), we still decompose the space $\mathbb{R}_{\xi}^{2}$ into three zones as in (3.4).
Case 2.1. The analysis in zone $A_{1}$
By [24, Section 10], one has
$(H_{\nu}^{\pm})^{\prime}(z)=H_{\nu-1}^{\pm}(z)-\frac{\nu}{z}H_{\nu}^{\pm},\quad J_{\nu}^{\prime}(z)=J_{\nu-1}(z)-\frac{\nu}{z}J_{\nu}(z).$ (3.36)
It follows from (2.6), (3.36) and direct computation that for $j=1,2$,
$|\partial_{\xi_{j}}\Psi_{0}(t,1,\xi)|\lesssim|\frac{\xi_{j}}{|\xi|}t^{\rho}H_{\rho-1}^{-}(|\xi|)H_{\rho}^{+}(t|\xi|)+\xi_{j}t^{\rho}\big{(}H_{\rho-2}^{-}(|\xi|)-\frac{\rho-1}{|\xi|}H_{\rho-1}^{-}(|\xi|)\big{)}H_{\rho}^{+}(t|\xi|)$ (3.37)
$\quad+\xi_{j}t^{\rho+1}\big{(}H_{\rho-1}^{+}(t|\xi|)-\frac{\rho}{t|\xi|}H_{\rho}^{+}(t|\xi|)\big{)}H_{\rho-1}^{-}(|\xi|)-\frac{\xi_{j}}{|\xi|}t^{\rho}H_{\rho-1}^{+}(|\xi|)H_{\rho}^{-}(t|\xi|)$
$\quad-\xi_{j}t^{\rho}\big{(}H_{\rho-2}^{+}(|\xi|)-\frac{\rho-1}{|\xi|}H_{\rho-1}^{+}(|\xi|)\big{)}H_{\rho}^{-}(t|\xi|)-\xi_{j}t^{\rho+1}\big{(}H_{\rho-1}^{-}(t|\xi|)-\frac{\rho}{t|\xi|}H_{\rho}^{-}(t|\xi|)\big{)}H_{\rho-1}^{+}(|\xi|)|$
$\lesssim\frac{|\xi_{j}|}{|\xi|}t^{\rho}H_{\rho-1}^{-}(|\xi|)H_{\rho}^{+}(t|\xi|)+|\xi_{j}|t^{\rho}H_{\rho-2}^{-}(|\xi|)H_{\rho}^{+}(t|\xi|)+|\xi_{j}|t^{\rho+1}H_{\rho-1}^{+}(t|\xi|)H_{\rho-1}^{-}(|\xi|)$
$\quad+\frac{|\xi_{j}|}{|\xi|}t^{\rho}H_{\rho-1}^{+}(|\xi|)H_{\rho}^{-}(t|\xi|)+|\xi_{j}|t^{\rho}H_{\rho-2}^{+}(|\xi|)H_{\rho}^{-}(t|\xi|)+|\xi_{j}|t^{\rho+1}H_{\rho-1}^{-}(t|\xi|)H_{\rho-1}^{+}(|\xi|)$
and
$|\partial_{\xi_{j}}\Psi_{1}(t,1,\xi)|\lesssim\frac{|\xi_{j}|}{|\xi|^{2}}t^{\rho}H_{\rho}^{-}(|\xi|)H_{\rho}^{+}(t|\xi|)+\frac{|\xi_{j}|}{|\xi|}t^{\rho}H_{\rho-1}^{-}(|\xi|)H_{\rho}^{+}(t|\xi|)+\frac{|\xi_{j}|}{|\xi|}t^{\rho+1}H_{\rho-1}^{+}(t|\xi|)H_{\rho}^{-}(|\xi|)$ (3.38)
$\quad+\frac{|\xi_{j}|}{|\xi|^{2}}t^{\rho}H_{\rho}^{+}(|\xi|)H_{\rho}^{-}(t|\xi|)+\frac{|\xi_{j}|}{|\xi|}t^{\rho}H_{\rho-1}^{+}(|\xi|)H_{\rho}^{-}(t|\xi|)+\frac{|\xi_{j}|}{|\xi|}t^{\rho+1}H_{\rho-1}^{-}(t|\xi|)H_{\rho}^{+}(|\xi|).$
Therefore, as in (3.17), it follows from (3.37)-(3.38), (2.3) and (2.22) that for $32<\mu<$ 3,
$\big{\|}\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\big{\|}_{L^{2}(A_{1})}\lesssim\big{\|}t^{-\frac{\mu}{2}}\hat{v}_{0}(\xi)+t^{-\frac{\mu}{2}}t(|\xi_{1}|+|\xi_{2}|)\hat{v}_{0}(\xi)\big{\|}_{L^{2}}+\big{\|}t^{-\frac{\mu}{2}}(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})\hat{v}_{0}(\xi)\big{\|}_{L^{2}}$
$\quad+\big{\|}t^{-\frac{\mu}{2}}(1+|\xi|^{2})^{-\frac{1}{2}}t(|\xi_{1}|+|\xi_{2}|)\hat{v}_{1}(\xi)\big{\|}_{L^{2}}+\big{\|}t^{-\frac{\mu}{2}}(1+|\xi|^{2})^{-\frac{1}{2}}(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})\hat{v}_{1}(\xi)\big{\|}_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\left\|t\partial_{1}v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\left\|t\partial_{2}v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\left\|(x_{1}\partial_{1}+x_{2}\partial_{2})v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}+2-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{1}v_{1}\right\|_{(1+\varepsilon_{1},2)}$ (3.39)
$\quad+t^{-\frac{\mu}{2}+2-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{2}v_{1}\right\|_{(1+\varepsilon_{1},2)}+t^{-\frac{\mu}{2}+2-\frac{2}{1+\varepsilon_{1}}}\left\|(x_{1}\partial_{1}+x_{2}\partial_{2})v_{1}\right\|_{(1+\varepsilon_{1},2)}+t^{-\frac{\mu}{2}+2-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{(1+\varepsilon_{1},2)}$
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}+2-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$
Case 2.2. The analysis in zone $A_{2}$
By (2.8) and (3.36), as treated in (3.37)-(3.38), we obtain that for $j=1,2$,
$|\partial_{\xi_{j}}\Psi_{0}(t,1,\xi)|\lesssim\frac{|\xi_{j}|}{|\xi|}t^{\rho}J_{-\rho+1}(|\xi|)J_{\rho}(t|\xi|)+\frac{|\xi_{j}|}{|\xi|}t^{\rho}J_{\rho-1}(|\xi|)J_{-\rho}(t|\xi|)+|\xi_{j}|t^{\rho}J_{-\rho}(|\xi|)J_{\rho}(t|\xi|)$ (3.40)
$\quad+t^{\rho}\frac{|\xi_{j}|}{|\xi|}J_{-\rho+1}(|\xi|)J_{\rho}(t|\xi|)+t^{\rho+1}|\xi_{j}|J_{\rho-1}(t|\xi|)J_{-\rho+1}(|\xi|)+t^{\rho}|\xi_{j}|J_{\rho-2}(|\xi|)J_{-\rho}(t|\xi|)$
$\quad+t^{\rho+1}|\xi_{j}|J_{-\rho-1}(t|\xi|)J_{\rho-1}(|\xi|)+t^{\rho}\frac{|\xi_{j}|}{|\xi|}J_{-\rho}(t|\xi|)J_{\rho-1}(|\xi|)$
and
$|\partial_{\xi_{j}}\Psi_{1}(t,1,\xi)|\lesssim\frac{|\xi_{j}|}{|\xi|}t^{\rho}J_{-\rho-1}(|\xi|)J_{\rho}(t|\xi|)+\frac{|\xi_{j}|}{|\xi|}t^{\rho+1}J_{\rho-1}(t|\xi|)J_{-\rho}(|\xi|)+\frac{|\xi_{j}|}{|\xi|}t^{\rho}J_{\rho-1}(|\xi|)J_{-\rho}(t|\xi|)$ (3.41)
$\quad+t^{\rho+1}\frac{|\xi_{j}|}{|\xi|}J_{-\rho-1}(t|\xi|)J_{\rho}(|\xi|)+t^{\rho}\frac{|\xi_{j}|}{|\xi|^{2}}J_{\rho}(t|\xi|)J_{-\rho}(|\xi|)+t^{\rho}\frac{|\xi_{j}|}{|\xi|^{2}}J_{-\rho}(|\xi|)J_{\rho}(t|\xi|)$
$\quad+t^{\rho}\frac{|\xi_{j}|}{|\xi|^{2}}J_{-\rho}(t|\xi|)J_{\rho}(|\xi|)+t^{\rho}\frac{|\xi_{j}|}{|\xi|^{2}}J_{-\rho}(t|\xi|)J_{\rho}(|\xi|).$
It follows from (3.40)-(3.41) and (2.24)-(2.25) that for $32<\mu<$ 3,
$\left\|\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\right\|_{L^{2}(A_{2})}\lesssim\big{\|}t^{-\frac{\mu}{2}}\hat{v}_{0}(\xi)\big{\|}_{L^{2}}+\big{\|}(1+t|\xi|)^{-\frac{\mu}{2}}t(|\xi_{1}|+|\xi_{2}|)(\hat{v}_{0}(\xi)+\hat{v}_{1}(\xi))\big{\|}_{L^{2}}$ (3.42)
$\qquad+\big{\|}(1+t|\xi|)^{-\frac{\mu}{2}}(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})(\hat{v}_{0}(\xi)+\hat{v}_{1}(\xi))\big{\|}_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{L^{2}}+t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{1}(v_{0}+v_{1})\right\|_{(1+\varepsilon_{1},2)}+t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{2}(v_{0}+v_{1})\right\|_{(1+\varepsilon_{1},2)}$
$\qquad+t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|(x_{1}\partial_{1}+x_{2}\partial_{2})(v_{0}+v_{1})\right\|_{(1+\varepsilon_{1},2)}+t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{0}+v_{1}\right\|_{(1+\varepsilon_{1},2)}$
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{L^{2}}+t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$
Case 2.3. The analysis in zone $A_{3}$
As in the zone $A_{2}$, by (3.40)-(3.41) and (2.25), one can obtain that for $t|\xi|\leq 1$,
$\left\|\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\right\|_{L^{2}(A_{3})}$ $\lesssim\big{\|}t^{-\frac{\mu}{2}}\hat{v}_{0}(\xi)\big{\|}_{L^{2}}+\big{\|}(1+t|\xi|)(1+t|\xi|)^{-1}(\hat{v}_{0}(\xi)+\hat{v}_{1}(\xi))\big{\|}_{L^{2}}$
$+\big{\|}(1+t|\xi|)(1+t|\xi|)^{-1}(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})(\hat{v}_{0}(\xi)+\hat{v}_{1}(\xi))\big{\|}_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\big{\|}v_{0}\big{\|}_{L^{2}}+t^{1-\frac{2}{1+\varepsilon_{1}}}\big{\|}v_{0}+v_{1}\big{\|}_{(1+\varepsilon_{1},2)}$ (3.43)
$\quad+t^{1-\frac{2}{1+\varepsilon_{1}}}\big{\|}(x_{1}\partial_{1}+x_{2}\partial_{2})(v_{0}+v_{1})\big{\|}_{(1+\varepsilon_{1},2)}$
$\lesssim t^{-\frac{\mu}{2}}\big{\|}v_{0}\big{\|}_{L^{2}}+t^{1-\frac{2}{1+\varepsilon_{1}}}\big{\|}v_{0}+v_{1}\big{\|}_{Z,1,(1+\varepsilon_{1},2)}.$
Therefore, combining (3.39), (3.42) and (3.43), the assertion (3.35) and further (3.29) are proved.
Case 3. The treatment of $\|L_{j}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}(j=1,2)$
We assert
$\|L_{j}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.44)
To prove (3.44), we only need to deal with $\|L_{1}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$ since the treatment on $\|L_{2}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$ is completely analogous. It follows from $t\geq 1$ and the support condition of the initial data that $|x|\lesssim t$ holds for $(t,x)\in\operatorname{supp}v$. Hence,
$\|L_{1}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\leq\|t\partial_{1}v\|_{L^{2}(\mathbb{R}^{2})}+\|x_{1}\partial_{t}v\|_{L^{2}(\mathbb{R}^{2})}\lesssim\|t|\xi_{1}|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}.$ (3.45)
Note that $\|t\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ can be estimated as in (3.34), it suffices only to treat $\|t|\xi_{1}|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$.
It follows from (2.6) and (2.22) that for $|\xi|\geq 1$,
$\left|\Psi_{0}(t,1,\xi)\right|\lesssim t^{-\frac{\mu}{2}},$
$\left|\Psi_{1}(t,1,\xi)\right|\lesssim t^{-\frac{\mu}{2}}|\xi|^{-1}.$
Then
$\left\|t|\xi_{1}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}$ $\lesssim t^{-\frac{\mu}{2}}\left\|t|\xi_{1}|\hat{v}_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\|(1+|\xi|^{2})^{-\frac{1}{2}}t|\xi_{1}|\hat{v}_{1}\|_{L^{2}}$ (3.46)
$\lesssim t^{-\frac{\mu}{2}}\left\|t\partial_{1}v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}+2-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{1}v_{1}\right\|_{(1+\varepsilon_{1},2)}$
$\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$
On the other hand, for $|\xi|\leq 1$ and $t|\xi|\geq 1$, it follows from (2.8), (2.24) and (2.25) that for $32<\mu<$ 3,
$\left|\Psi_{0}(t,1,\xi)\right|\lesssim t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}}\lesssim(1+t|\xi|)^{-1},$
$\left|\Psi_{1}(t,1,\xi)\right|\lesssim t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}}\lesssim(1+t|\xi|)^{-1}.$
Then
$\left\|t|\xi_{1}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}$ $\lesssim\|(1+t|\xi|)^{-1}t|\xi_{1}|(\hat{v}_{0}+\hat{v}_{1})\|_{L^{2}}$ (3.47)
$\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{1}(v_{0}+v_{1})\right\|_{(1+\varepsilon_{1},2)}$
$\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|(v_{0}+v_{1})\right\|_{Z,1,(1+\varepsilon_{1},2)}.$
For $t|\xi|\leq 1$, by (2.8) and (2.24), we have
$\left|\Psi_{0}(t,1,\xi)\right|\lesssim 1=(1+t|\xi|)(1+t|\xi|)^{-1}\lesssim(1+t|\xi|)^{-1},$
$\left|\Psi_{1}(t,1,\xi)\right|\lesssim 1=(1+t|\xi|)(1+t|\xi|)^{-1}\lesssim(1+t|\xi|)^{-1}.$
Then
$\left\|t|\xi_{1}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{3})}\lesssim\|(1+t|\xi|)^{-1}(\hat{v}_{0}+\hat{v}_{1})\|_{L^{2}}\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{0}+v_{1}\right\|_{(1+\varepsilon_{1},2)}.$ (3.48)
Therefore, combining (3.46), (3.47) and (3.48) gives
$\|t|\xi_{1}|\hat{v}(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.49)
Together with (3.34), this yields (3.44) immediately.
Case 4. The treatment of $\|\Omega_{12}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$
We now start to prove
$\|\Omega_{12}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$ (3.50)
As treated in (3.45), one has
$\|\Omega_{12}v(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$ $\leq\|\frac{x_{1}}{t}t\partial_{2}v\|_{L^{2}(\mathbb{R}^{2})}+\|\frac{x_{2}}{t}t\partial_{1}v\|_{L^{2}(\mathbb{R}^{2})}$ (3.51)
$\lesssim\|t|\xi_{2}|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t|\xi_{1}|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}.$
Note that by (3.49), it holds that for $j=1,2$,
$\|t|\xi_{j}|\hat{v}(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|v_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\left\|v_{0}+v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$
Therefore, (3.50) is shown.
Finally, collecting the results in Cases 1-4, Lemma 3.1 is proved. ∎
Based on some crucial estimates in Lemma 3.1, we now treat the first order derivative of solution $v$ to (3.1).
Lemma 3.2.
Let $v$ be the solution to (3.1). Then for $32<\mu<$ 3,
$\|\partial v(t,\cdot)\|_{Z,1,2}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$ (3.52)
Proof.
By the definition of $\|\partial v\|_{Z,1,2}$, we firstly prove
$\|\partial v(t,\cdot)\|_{L^{2}}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{L^{2}}+t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}.$ (3.53)
In fact, as in Lemma 3.1, we decompose the frequency space $\mathbb{R}_{\xi}^{2}$ into three zones $A_{1}$, $A_{2}$ and $A_{3}$. For $|\xi|\geq 1$, it follows from (2.6), (3.5) and Parseval’s equality that
$\left\||\xi|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}$ $\leq\left\|\Psi_{0}(t,1,\xi)\right\|_{L^{\infty}}\left\||\xi|\hat{v}_{0}\right\|_{L^{2}}+\left\||\Psi_{1}(t,1,\xi)||\xi|\hat{v}_{1}\right\|_{L^{2}}$ (3.54)
$\lesssim t^{-\frac{\mu}{2}}\left\||\xi|\hat{v}_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\|\hat{v}_{1}\|_{L^{2}}$
$=t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\|v_{1}\|_{L^{2}}.$
For $|\xi|\leq 1$ and $t|\xi|\geq 1$, by (2.8) and (3.10), one has
$\left\||\xi|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}\lesssim\big{\|}(t|\xi|)^{-\frac{\mu}{2}}|\xi|(\hat{v}_{0}+\hat{v}_{1})\big{\|}_{L^{2}}\lesssim t^{-1}\big{\|}(t|\xi|)^{1-\frac{\mu}{2}}(\hat{v}_{0}+\hat{v}_{1})\big{\|}_{L^{2}}\lesssim t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}.$ (3.55)
For $t|\xi|\leq 1$, by (2.8) and (3.13), we arrive at
$\left\||\xi|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{3})}\lesssim\left\||\xi|(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}\leq t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}.$
This, together with (3.54) and (3.55), yields
$\left\||\xi|\hat{v}(t,\cdot)\right\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{L^{2}}+t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}.$
Analogously, by (3.16), (3.20) and (3.24), we have
$\left\|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{L^{2}}+t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}.$
Thus, (3.53) is shown.
Next we turn to the proof of (3.52). As treated as in Lemma 3.1, the proof procedure is divided into the following four cases.
Case 1. The treatment of $\|\partial(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$
We now show
$\|\partial(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$ (3.56)
Note that
$\|\partial(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$ $\lesssim\|\partial_{t}^{2}v\|_{L^{2}(\mathbb{R}^{2})}+\sum_{j=1}^{2}\|\partial_{t}\partial_{j}v\|_{L^{2}(\mathbb{R}^{2})}+\sum_{i,j=1}^{2}\|\partial_{i}\partial_{j}v\|_{L^{2}(\mathbb{R}^{2})}$ (3.57)
$\lesssim\|\partial_{t}^{2}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\left\||\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(\mathbb{R}^{2})}+\sum_{i,j=1}^{2}\left\||\xi_{j}||\xi_{i}|\hat{v}(t,\cdot)\right\|_{L^{2}(\mathbb{R}^{2})}.$
It follows from Lemma 2.1 and Lemma 2.2 that for $t|\xi|\leq 1$,
$\left|\partial_{t}^{2}\Psi_{0}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{3}(t|\xi|)^{-\frac{1}{2}}|\xi|^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|\cdot|\xi|,$ (3.58)
$\left|\partial_{t}^{2}\Psi_{1}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{2}(t|\xi|)^{-\frac{1}{2}}|\xi|^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}\cdot|\xi|.$
For $|\xi|\leq 1$ and $t|\xi|\geq 1$,
$\left|\partial_{t}^{2}\Psi_{0}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{3}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho-1}=(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|,$ (3.59)
$\left|\partial_{t}\Psi_{1}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{2}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho}=(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|.$
For $t|\xi|\leq 1$,
$\left|\partial_{t}^{2}\Psi_{0}(t,1,\xi)\right|\lesssim t^{-1}|\xi|,$ (3.60)
$\left|\partial_{t}\Psi_{1}(t,1,\xi)\right|\lesssim t^{-1}|\xi|.$
By comparing the estimates in (3.58), (3.59) and (3.60) with the corresponding ones in (3.16), (3.20) and (3.24), it is known that the treatment of $\|\partial_{t}^{2}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ can be similar to the one for $\left\||\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(\mathbb{R}^{2})}$. Therefore, in order to prove (3.56), it suffices only to deal with $\left\||\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(\mathbb{R}^{2})}$ and $\left\||\xi_{j}||\xi_{i}|\hat{v}(t,\cdot)\right\|_{L^{2}(\mathbb{R}^{2})}$ $(i,j=1,2)$.
It follows from (3.16) and (3.5) that for $|\xi|\geq 1$,
$\left\||\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}$ $\lesssim t^{-\frac{\mu}{2}}\left\||\xi||\xi|\hat{v_{0}}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\||\xi|\hat{v}_{1}\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|(|\xi_{1}|+|\xi_{2}|)\hat{v}_{1}\right\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}$
and
$\left\||\xi_{j}||\xi_{i}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}$ $\lesssim t^{-\frac{\mu}{2}}\left\||\xi||\xi_{j}|v_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\||\xi_{j}|\hat{v}_{1}\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}.$
For $|\xi|\leq 1$ and $t|\xi|\geq 1$, by (3.10) and (3.20), one has that for $j=1,2$,
$\left\||\xi_{j}||\xi_{i}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}\lesssim\big{\|}(t|\xi|)^{-\frac{\mu}{2}}|\xi|(\hat{v}_{0}+\hat{v}_{1})\big{\|}_{L^{2}}\lesssim t^{-1}\big{\|}(t|\xi|)^{1-\frac{\mu}{2}}(\hat{v}_{0}+\hat{v}_{1})\big{\|}_{L^{2}}\lesssim t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}$
and
$\left\||\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}\lesssim\big{\|}(t|\xi|)^{-\frac{\mu}{2}}|\xi|(\hat{v}_{0}+\hat{v}_{1})\big{\|}_{L^{2}}\lesssim t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}.$
For $t|\xi|\leq 1$, by (3.13) and (3.24), we get
$\left\||\xi_{j}||\xi_{i}|\hat{v}(t,\cdot)\right\|_{L^{2}(A_{3})}\lesssim\left\||\xi|^{2}(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}\lesssim t^{-2}\left\|v_{0}+v_{1}\right\|_{L^{2}}$
and
$\left\||\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{3})}\lesssim\left\|t^{-1}|\xi|(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}\lesssim t^{-2}\left\|v_{0}+v_{1}\right\|_{L^{2}}.$
Collecting the above conclusions, (3.56) is then proved.
Case 2. The treatment of $\|L_{0}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$
We now prove
$\|L_{0}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$ (3.61)
Analogous to the treatment of (3.57) and (3.30), it suffices to estimate $\|t|\xi|\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ and $\||\xi|(\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}})\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$.
At first, we estimate $\|t|\xi|\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$. For $|\xi|\geq 1$, by (3.16), one has
$\left\|t|\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{1})}\lesssim t^{-\frac{\mu}{2}}\left\|t|\xi||\xi|\hat{v}_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\|t|\xi|\hat{v}_{1}\|_{L^{2}}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}.$ (3.62)
For $|\xi|\leq 1$ and $t|\xi|\geq 1$, by (3.20), we arrive at
$\left\|t|\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{2})}$ $\lesssim t^{-1}\big{\|}(t|\xi|)^{1-\frac{\mu}{2}}t|\xi|(\hat{v}_{0}+\hat{v}_{1})\big{\|}_{L^{2}}$ (3.63)
$\lesssim t^{-1}\left\|(t\partial_{1}+x_{1}\partial_{t})(v_{0}+v_{1})\right\|_{L^{2}}+t^{-1}\left\|(t\partial_{1}+x_{1}\partial_{t})(v_{0}+v_{1})\right\|_{L^{2}}$
$\lesssim t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$
For $t|\xi|\leq 1$, by (3.24), one can obtain
$\left\|t|\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(A_{3})}\leq\big{\|}t^{-1}t|\xi|(\hat{v}_{0}+\hat{v}_{1})\big{\|}_{L^{2}}\lesssim t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}.$
This, together with (3.62) and (3.63), yields
$\|t|\xi|\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$ (3.64)
Next we estimate $\||\xi|(\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}})\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$.
For $|\xi|\geq 1$, it follows from (3.36)-(3.37) and (2.22) that
$\left\||\xi|(\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}})\hat{v}\right\|_{L^{2}(A_{1})}$ (3.65)
$=\||\xi|\xi_{1}(\partial_{\xi_{1}}\Psi_{0})\hat{v}_{0}(\xi)+|\xi|\xi_{1}\Psi_{0}\partial_{\xi_{1}}\hat{v}_{0}(\xi)+|\xi|\xi_{1}(\partial_{\xi_{1}}\Psi_{1})\hat{v}_{1}(\xi)+|\xi|\xi_{1}\Psi_{1}\partial_{\xi_{1}}\hat{v}_{1}(\xi)$
$\quad+|\xi|\xi_{2}(\partial_{\xi_{2}}\Psi_{0})\hat{v}_{0}(\xi)+|\xi|\xi_{2}\Psi_{0}\partial_{\xi_{2}}\hat{v}_{0}(\xi)+|\xi|\xi_{2}(\partial_{\xi_{2}}\Psi_{1})\hat{v}_{1}(\xi)+|\xi|\xi_{2}\Psi_{1}\partial_{\xi_{2}}\hat{v}_{1}(\xi)\|_{L^{2}}$
$\lesssim\big{\|}t^{-\frac{\mu}{2}}|\xi|\hat{v}_{0}(\xi)+t^{-\frac{\mu}{2}}t(|\xi_{1}|+|\xi_{2}|)|\xi|\hat{v}_{0}(\xi)+t^{-\frac{\mu}{2}}|\xi|(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})\hat{v}_{0}(\xi)\big{\|}_{L^{2}}$
$\quad+\big{\|}t^{-\frac{\mu}{2}}t(|\xi_{1}|+|\xi_{2}|)\hat{v}_{1}(\xi)\big{\|}_{L^{2}}+\big{\|}t^{-\frac{\mu}{2}}(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})\hat{v}_{1}(\xi)\big{\|}_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|t\partial_{1}v_{1}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\left\|t\partial_{2}v_{1}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\left\|(x_{1}\partial_{1}+x_{2}\partial_{2})v_{1}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}.$
For $|\xi|\leq 1$ and $t|\xi|\geq 1$, by (3.40)-(3.41) and (2.24)-(2.25), one has that
$\left\||\xi|(\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}})\hat{v}\right\|_{L^{2}(A_{2})}\lesssim\big{\|}t^{-\frac{\mu}{2}}|\xi|\hat{v}_{0}(\xi)+t^{-\frac{\mu}{2}}t(|\xi_{1}|+|\xi_{2}|)|\xi|\hat{v}_{0}(\xi)\big{\|}_{L^{2}}$ (3.66)
$\quad+t^{-1}\big{\|}(t|\xi|)^{1-\frac{\mu}{2}}t(|\xi_{1}|+|\xi_{2}|)(\hat{v}_{0}(\xi)+\hat{v}_{1}(\xi))\big{\|}_{L^{2}}+t^{-1}\big{\|}(t|\xi|)^{1-\frac{\mu}{2}}(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})(\hat{v}_{0}(\xi)+\hat{v}_{1}(\xi))\big{\|}_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-1}\left\|(t\partial_{1}+x_{1}\partial_{t})(v_{0}+v_{1})\right\|_{L^{2}}+t^{-1}\left\|(t\partial_{2}+x_{2}\partial_{t})(v_{0}+v_{1})\right\|_{L^{2}}$
$\quad+t^{-1}\left\|(x_{1}\partial_{1}+x_{2}\partial_{2}+t\partial_{t})(v_{0}+v_{1})\right\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$
For $t|\xi|\leq 1$, we have from (3.40)-(3.41) and (2.25) that
$\left\||\xi|(\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}})\hat{v}\right\|_{L^{2}(A_{3})}$ (3.67)
$\lesssim\big{\|}t^{-\frac{\mu}{2}}|\xi|\hat{v}_{0}(\xi)\big{\|}_{L^{2}}+\big{\|}|\xi|(\hat{v}_{0}(\xi)+\hat{v}_{1}(\xi))\big{\|}_{L^{2}}+\left\||\xi|(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})(\hat{v}_{0}(\xi)+\hat{v}_{1}(\xi))\right\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}+t^{-1}\left\|(x_{1}\partial_{1}+x_{2}\partial_{2})(v_{0}+v_{1})\right\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$
Therefore, (3.61) follows from (3.64) and (3.65)-(3.67).
Case 3. The treatment of $\|L_{j}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}(j=1,2)$
We now establish
$\|L_{j}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$ (3.68)
It suffices only to estimate $\|L_{1}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$. Note that $|x|\lesssim t$ holds for $(t,x)\in\operatorname{supp}v$. Then
$\|L_{1}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$ $\leq\|t|\xi_{1}||\xi|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t|\xi_{1}|\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|\frac{x_{1}}{t}t\partial_{t}\partial v\|_{L^{2}(\mathbb{R}^{2})}$ (3.69)
$\lesssim\|t|\xi_{1}||\xi|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t|\xi_{1}|\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t\partial_{t}\partial v\|_{L^{2}(\mathbb{R}^{2})}.$
We next deal with each term in the right hand side of (3.69).
By comparing the estimates of (3.5), (3.10) and (3.13) with those of (3.16), (3.20) and (3.24), respectively, then $\|t|\xi_{1}||\xi|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ can be treated as for $\|t|\xi_{1}|\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ in (3.64).
For the term $\|t\partial_{t}\partial v\|_{L^{2}(\mathbb{R}^{2})}$, the following estimate has been obtained in the proof procedure of $\|L_{0}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$ in Case 2.
$\|t\partial_{t}\partial v\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$
Finally, we estimate the term $\|t|\xi_{1}|\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$. Analogously to the treatment in (3.53), one has that for $|\xi|\geq 1$,
$\left\|t|\xi_{1}|\partial_{t}\hat{v}\right\|_{L^{2}(A_{1})}$ $\leq\left\|\partial_{t}\Psi_{0}(t,1,\xi)t|\xi_{1}|\hat{v}_{1}\right\|_{L^{2}}+\left\|\partial_{t}\Psi_{1}(t,1,\xi)t|\xi_{1}|\hat{v}_{1}\right\|_{L^{2}}$ (3.70)
$\lesssim t^{-\frac{\mu}{2}}\left\|t|\xi_{1}||\xi|\hat{v}_{0}\right\|_{L^{2}}+t^{-\frac{\mu}{2}}\|t|\xi_{1}|\hat{v}_{1}\|_{L^{2}}$
$\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\|v_{1}\|_{Z,1,2}.$
In addition, for $|\xi|\leq 1$ and $t|\xi|\geq 1$, it holds
$\left\|t|\xi_{1}|\partial_{t}\hat{v}\right\|_{L^{2}(A_{2})}\lesssim t^{-1}\big{\|}(t|\xi|)^{1-\frac{\mu}{2}}t|\xi_{1}|(\hat{v}_{0}+\hat{v}_{1})\big{\|}_{L^{2}}\lesssim t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$ (3.71)
For $t|\xi|\leq 1$,
$\left\|t|\xi_{1}|\partial_{t}\hat{v}\right\|_{L^{2}(A_{3})}\lesssim\left\||\xi|(\hat{v}_{0}+\hat{v}_{1})\right\|_{L^{2}}\leq t^{-1}\left\|v_{0}+v_{1}\right\|_{L^{2}}.$
Therefore,
$\left\|t|\xi_{1}|\partial_{t}\hat{v}\right\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2},$ (3.72)
which derives (3.68).
Case 4. The treatment of $\|\Omega_{12}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$
We will prove
$\|\Omega_{12}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$ (3.73)
Analogously treated as in (3.51) and (3.69), one can obtain
$\|\Omega_{12}(\partial v)(t,\cdot)\|_{L^{2}(\mathbb{R}^{2})}$ $\leq\|\frac{x_{1}}{t}t\partial_{2}\partial v\|_{L^{2}(\mathbb{R}^{2})}+\|\frac{x_{2}}{t}t\partial_{1}v\|_{L^{2}(\mathbb{R}^{2})}$ (3.74)
$\lesssim\|t|\xi_{2}||\xi|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t|\xi_{1}||\xi|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$
$\quad+\left\|t|\xi_{2}|\partial_{t}\hat{v}\right\|_{L^{2}(\mathbb{R}^{2})}+\left\|t|\xi_{1}|\partial_{t}\hat{v}\right\|_{L^{2}(\mathbb{R}^{2})}.$
In addition, by (3.72), we have that for $j=1,2$,
$\left\|t|\xi_{j}|\partial_{t}\hat{v}\right\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$
On the other hand, as illustrated in (3.69), it holds that for $j=1,2$,
$\left\|t|\xi_{j}||\xi|\hat{v}\right\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{-\frac{\mu}{2}}\left\|\nabla v_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\left\|v_{1}\right\|_{Z,1,2}+t^{-1}\left\|v_{0}+v_{1}\right\|_{Z,1,2}.$
Therefore, (3.73) is shown.
Collecting all the cases above, we complete the proof of Lemma 3.2.
4 Estimates of solution to 2-D inhomogeneous equation $\square w+\frac{\mu}{t}\,\partial_{t}w=F$
In this section, we will derive some time-decay estimates of solution to the 2-D linear problem
[…]
where $32<\mu<$ 3. To this end, it follows from the Duhamel’s principle that we first treat such a homogeneous initial data problem from the initial variable time $\tau\geq 1$
[…]
where $v_{1}(x)\in C_{0}^{\infty}(\mathbb{R}^{2})$.
Lemma 4.1.
Let $v$ solve (4.2). Then for any $\varepsilon_{1}\in(0,1)$, there exists a constant $\delta(\varepsilon_{1})=\frac{2\varepsilon_{1}}{1+\varepsilon_{1}}>0$ such that
$\|v\|_{Z,1,2}\lesssim t^{\delta(\varepsilon_{1})-1}\tau\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}$ (4.3)
and
$\|\partial v\|_{Z,1,2}\lesssim t^{-1}\tau\left\|v_{1}\right\|_{Z,1,2}.$ (4.4)
Proof.
Although the proof procedure of Lemma 4.1 is similar to that for Lemma 3.1 and Lemma 3.2, we still provide some details of proof for (4.3) and (4.4) due to the appearance of the parameter $\tau$.
We now establish
$\|v\|_{Z,1,2}\lesssim t^{\delta(\varepsilon_{1})-1}\tau\left\|v_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}.$
Note that
$\|v\|_{Z,1,2}\leq\|\hat{v}\|_{L^{2}}+\sum_{j=1}^{2}\||\xi_{j}|\hat{v}\|_{L^{2}}+\|\partial_{t}\hat{v}\|_{L^{2}}+\|\widehat{L_{0}v}\|_{L^{2}}+\sum_{j=1}^{2}\|\widehat{L_{j}v}\|_{L^{2}}+\|\widehat{\Omega_{12}v}\|_{L^{2}}.$ (4.5)
Next we deal with each term in (4.5).
It follows from (2.6) and (2.22) that for $t|\xi|\geq\tau|\xi|\geq 1$,
$\left|\Psi_{1}(t,\tau,\xi)\right|\lesssim t^{\rho}\tau^{1-\rho}(\tau|\xi|)^{-\frac{1}{2}}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}|\xi|^{-1}.$ (4.6)
In addition, for $t|\xi|\geq 1$ and $\tau|\xi|\leq 1$, by (2.8) and (2.24)-(2.25), one has
$\left|\Psi_{1}(t,\tau,\xi)\right|\lesssim t^{\rho}\tau^{1-\rho}(\tau|\xi|)^{\rho}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}}\tau.$ (4.7)
For $\tau|\xi|\leq t|\xi|\leq 1$, by (2.8) and (2.25), we have
$\left|\Psi_{1}(t,\tau,\xi)\right|$ $\lesssim t^{\rho}\tau^{1-\rho}\left[(\tau|\xi|)^{-\rho}(t|\xi|)^{\rho}+(\tau|\xi|)^{\rho}(t|\xi|)^{-\rho}\right]$ (4.8)
$=t^{2\rho}\tau^{1-2\rho}+\tau\lesssim\tau.$
Thus it follows from (4.6)-(4.8) and (2.3) with $\hat{v}_{0}=0$ that for $t\geq\tau\geq 1$,
$\|\hat{v}\|_{L^{2}}$ $\lesssim\big{\|}t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}|\xi|^{-1}\hat{v}_{1}(\xi)\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}+\big{\|}t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}}\tau\hat{v}_{1}(\xi)\big{\|}_{L^{2}(\tau|\xi|\leqslant 1,t|\xi|\geqslant 1)}+\left\|\tau\hat{v}_{1}(\xi)\right\|_{L^{2}(t|\xi|\leqslant 1)}$ (4.9)
$\lesssim t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}\cdot\tau\left\|(1+\tau|\xi|)^{-1}\hat{v}_{1}\right\|_{L^{2}}+\tau\left\|(1+t|\xi|)^{-1}\hat{v}_{1}\right\|_{L^{2}}+\tau\left\|(1+\tau|\xi|)(1+\tau|\xi|)^{-1}\hat{v}_{1}\right\|_{L^{2}}$
$\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\tau\left\|v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}+\tau t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}+\tau\cdot\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}$
$\lesssim t^{\delta(\varepsilon_{1})-1}\tau\left\|v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}$
and
$\|t|\xi_{j}|\hat{v}\|_{L^{2}}$ $\lesssim\big{\|}t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}|\xi|^{-1}t|\xi_{j}|\hat{v}_{1}\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}+\big{\|}t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}}\tau t|\xi_{j}|\hat{v}_{1}\big{\|}_{L^{2}(\tau|\xi|\leqslant 1,t|\xi|\geqslant 1)}+\left\|\tau t|\xi_{j}|\hat{v}_{1}\right\|_{L^{2}(t|\xi|\leqslant 1)}$ (4.10)
$\lesssim t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}+1}\left\|(1+\tau|\xi|)^{-1}t|\xi_{j}|\hat{v}_{1}\right\|_{L^{2}}+\tau\left\|(1+t|\xi|)^{-1}t|\xi_{j}|\hat{v}_{1}\right\|_{L^{2}}+\tau\left\|(1+\tau|\xi|)^{-1}t|\xi_{j}|\hat{v}_{1}\right\|_{L^{2}}$
$\lesssim\tau\cdot\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}+\tau\cdot t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}+\tau\cdot\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}$
$\lesssim t^{\delta(\varepsilon_{1})-1}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}.$
Then
$\||\xi_{j}|\hat{v}\|_{L^{2}}=t^{-1}\|t|\xi_{j}|\partial_{t}\hat{v}\|_{L^{2}}\lesssim t^{-\frac{2}{1+\varepsilon_{1}}}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}\leq t^{\delta(\varepsilon_{1})-1}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}.$ (4.11)
In addition, for $t|\xi|\geq\tau|\xi|\geq 1$, we have from (2.6) and (2.22) that
$\left|\partial_{t}\Psi_{1}(t,\tau,\xi)\right|\lesssim|\xi|t^{\rho}\tau^{1-\rho}(\tau|\xi|)^{-\frac{1}{2}}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}.$ (4.12)
For $t|\xi|\geq 1$ and $\tau|\xi|\leq 1$, by (2.8) and (2.24)-(2.25), one has
$\left|\partial_{t}\Psi_{1}(t,\tau,\xi)\right|\lesssim|\xi|t^{\rho}\tau^{1-\rho}(\tau|\xi|)^{-\rho}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|^{\frac{\mu}{2}}\tau^{\mu}.$ (4.13)
For $\tau|\xi|\leq t|\xi|\leq 1$, by (2.8) and (2.25) it holds
$\left|\partial_{t}\Psi_{1}(t,\tau,\xi)\right|$ $\lesssim|\xi|t^{\rho}\tau^{1-\rho}\left[(\tau|\xi|)^{-\rho}(t|\xi|)^{\rho-1}+(\tau|\xi|)^{\rho}(t|\xi|)^{1-\rho}\right]$ (4.14)
$\leq t^{-1}\tau.$
Hence we have
$\|t\partial_{t}\hat{v}\|_{L^{2}}$ $\lesssim\big{\|}t^{1-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}\hat{v}_{1}(\xi)\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}+\big{\|}t^{1-\frac{\mu}{2}}|\xi|^{\frac{\mu}{2}}\tau^{\mu}\hat{v}_{1}(\xi)\big{\|}_{L^{2}(\tau|\xi|\leqslant 1,t|\xi|\geqslant 1)}+\left\|\tau\hat{v}_{1}(\xi)\right\|_{L^{2}(t|\xi|\leqslant 1)}$ (4.15)
$\lesssim t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}+1}\left\|(1+\tau|\xi|)^{-1}t|\xi|\hat{v}_{1}\right\|_{L^{2}}+t^{1-\frac{\mu}{2}}\tau^{\frac{\mu}{2}-1}\big{\|}\tau^{\frac{\mu}{2}+1}|\xi|^{\frac{\mu}{2}}(1+t|\xi|)^{-1}t|\xi|\hat{v}_{1}\big{\|}_{L^{2}(\tau|\xi|\leqslant 1,t|\xi|\geqslant 1)}$
$\quad+\tau\left\|(1+\tau|\xi|)(1+\tau|\xi|)^{-1}\hat{v}_{1}\right\|_{L^{2}}$
$\lesssim\tau\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{1}v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}+\tau\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{2}v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}+\tau t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{1}v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}$
$\quad+\tau t^{1-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{2}v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}+\tau\cdot\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}$
$\lesssim t^{\delta(\varepsilon_{1})-1}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}$
and
$\|\partial_{t}\hat{v}\|_{L^{2}}=t^{-1}\|t\partial_{t}\hat{v}\|_{L^{2}}\lesssim t^{\delta(\varepsilon_{1})-1}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}.$ (4.16)
By an analogous analysis in the proof process of (3.35), one has that for $j=1,2$, and $\tau|\xi|\geq 1$,
$\partial_{\xi_{j}}$ $\Psi_{1}(t,\tau,\xi)=\frac{i\pi}{4}t^{\rho}\tau^{1-\rho}\left[\frac{\tau\xi_{j}}{|\xi|}\big{(}H_{\rho-1}^{-}(\tau|\xi|)-\frac{\rho}{\tau|\xi|}H_{\rho}^{-}(\tau|\xi|)\big{)}H_{\rho}^{+}(t|\xi|)+\frac{t\xi_{j}}{|\xi|}\left(H_{\rho-1}^{+}(t|\xi|)\right.\right.$ (4.17)
$\quad-\left.\frac{\rho}{t|\xi|}H_{\rho}^{+}(t|\xi|)\right)H_{\rho}^{-}(\tau|\xi|)-\frac{\tau\xi_{j}}{|\xi|}\big{(}H_{\rho-1}^{+}(\tau|\xi|)-\frac{\rho}{\tau|\xi|}H_{\rho}^{+}(\tau|\xi|)\big{)}H_{\rho}^{-}(t|\xi|)$
$\quad\left.-\frac{t\xi_{j}}{|\xi|}\big{(}H_{\rho-1}^{-}(t|\xi|)-\frac{\rho}{t|\xi|}H_{\rho}^{-}(t|\xi|)\big{)}H_{\rho}^{+}(\tau|\xi|)\right]$
$=\frac{i\pi}{4}t^{\rho}\tau^{1-\rho}\left[\frac{\tau\xi_{j}}{|\xi|}H_{\rho-1}^{-}(\tau|\xi|)H_{\rho}^{+}(t|\xi|)+\frac{t\xi_{j}}{|\xi|}H_{\rho-1}^{+}(t|\xi|)H_{\rho}^{-}(\tau|\xi|)\right.$
$\quad-\frac{\tau\xi_{j}}{|\xi|}H_{\rho-1}^{+}(\tau|\xi|)H_{\rho}^{-}(t|\xi|)-\frac{t\xi_{j}}{|\xi|}H_{\rho-1}^{-}(t|\xi|)H_{\rho}^{+}(\tau|\xi|)-\frac{\xi_{j}\rho}{|\xi|^{2}}H_{\rho}^{-}(\tau|\xi|)H_{\rho}^{+}(t|\xi|)$
$\left.\quad+\frac{\xi_{j}\rho}{|\xi|^{2}}H_{\rho}^{+}(t|\xi|)H_{\rho}^{-}(\tau|\xi|)+\frac{\xi_{j}\rho}{|\xi|^{2}}H_{\rho}^{+}(\tau|\xi|)H_{\rho}^{-}(t|\xi|)+\frac{\xi_{j}\rho}{|\xi|^{2}}H_{\rho}^{-}(t|\xi|)H_{\rho}^{+}(\tau|\xi|)\right]$
and for $\tau|\xi|\leq 1$,
$\partial_{\xi_{j}}$ $\Psi_{1}(t,\tau,\xi)=\frac{\pi}{2}\csc(\rho\pi)t^{\rho}\tau^{1-\rho}\left[\frac{\tau\xi_{j}}{|\xi|}\big{(}J_{-\rho-1}(\tau|\xi|)+\frac{\rho}{\tau|\xi|}J_{-\rho}(\tau|\xi|)\big{)}J_{\rho}(t|\xi|)+\frac{t\xi_{j}}{|\xi|}\bigg{(}J_{\rho-1}(t|\xi|)\right.$ (4.18)
$\quad-\frac{\rho}{t|\xi|}J_{\rho}(t|\xi|)\bigg{)}J_{-\rho}(\tau|\xi|)-\frac{\tau\xi_{j}}{|\xi|}\big{(}J_{\rho-1}(\tau|\xi|)-\frac{\rho}{\tau|\xi|}J_{\rho}(\tau|\xi|)\big{)}J_{-\rho}(t|\xi|)$
$\quad\left.-\frac{t\xi_{j}}{|\xi|}\big{(}J_{-\rho-1}(t|\xi|)+\frac{\rho}{t|\xi|}J_{\rho}(t|\xi|)\big{)}J_{\rho}(\tau|\xi|)\right]$
$=\frac{\pi}{2}\csc(\rho\pi)t^{\rho}\tau^{1-\rho}\left[\frac{\tau\xi_{j}}{|\xi|}J_{-\rho-1}(\tau|\xi|)J_{\rho}(t|\xi|)+\frac{t\xi_{j}}{|\xi|}J_{\rho-1}(t|\xi|)J_{-\rho}(\tau|\xi|)\right.$
$\quad-\frac{\tau\xi_{j}}{|\xi|}J_{\rho-1}(\tau|\xi|)J_{-\rho}(t|\xi|)-\frac{t\xi_{j}}{|\xi|}J_{-\rho-1}(t|\xi|)J_{\rho}(\tau|\xi|)+\frac{\xi_{j}\rho}{|\xi|^{2}}J_{-\rho}(\tau|\xi|)J_{\rho}(t|\xi|)$
$\left.\quad-\frac{\xi_{j}\rho}{|\xi|^{2}}J_{\rho}(t|\xi|)J_{-\rho}(\tau|\xi|)+\frac{\xi_{j}\rho}{|\xi|^{2}}J_{\rho}(\tau|\xi|)J_{-\rho}(t|\xi|)-\frac{\xi_{j}\rho}{|\xi|^{2}}J_{-\rho}(t|\xi|)J_{\rho}(\tau|\xi|)\right].$
As in (3.39) with $\hat{v}_{0}=0$, it follows from (4.17) that
$\left\|\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\right\|_{L^{2}(\tau|\xi|\geq 1)}$ (4.19)
$=\|\xi_{1}(\partial_{\xi_{1}}\Psi_{1})\hat{v_{1}}(\xi)+\xi_{1}\Psi_{1}\partial_{\xi_{1}}\hat{v_{1}}(\xi)+\xi_{2}(\partial_{\xi_{2}}\Psi_{1})\hat{v_{1}}(\xi)+\xi_{2}\Psi_{1}\partial_{\xi_{2}}\hat{v_{1}}(\xi)\|_{L^{2}(\tau|\xi|\geq 1)}$
$\lesssim\big{\|}t^{1-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}\hat{v}_{1}(\xi)\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}+\big{\|}t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}|\xi|^{-1}\left(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}}\right)\hat{v}_{1}\left(\xi\right)\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}$
$\lesssim t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}+1}\left\|(1+\tau|\xi|)^{-1}t|\xi|\hat{v}_{1}\right\|_{L^{2}}+\tau\left\|(1+\tau|\xi|)^{-1}\left(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}}\right)\hat{v}_{1}\left(\xi\right)\right\|_{L^{2}}$
$\lesssim\tau\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{1}v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}+\tau\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|t\partial_{2}v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}+\tau\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|(x_{1}\partial_{1}+x_{2}\partial_{2})v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}$
$\quad+\tau\tau^{1-\frac{2}{1+\varepsilon_{1}}}\left\|v_{1}\right\|_{\left(1+\varepsilon_{1},2\right)}$
$\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}.$
Similarly, by (4.18), one can easily obtain
$\left\|\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\right\|_{L^{2}(\tau|\xi|\leq 1)}\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}.$ (4.20)
Therefore, by (4.9), (4.15) and (4.19)-(4.20), we arrive at
$\|\widehat{L_{0}v}\|_{L^{2}}\lesssim 2\|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}.$
Meanwhile, by (4.11) and (4.16), one has
$\|\widehat{\partial v}\|_{L^{2}}\lesssim\sum_{j=1}^{2}\||\xi_{j}|\hat{v}\|_{L^{2}}+\|\partial_{t}\hat{v}\|_{L^{2}}\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}.$
In addition, as in (3.45) and (3.51), it follows from (4.10) and (4.15) that
$\sum_{j=1}^{2}\|\widehat{L_{j}v}\|_{L^{2}}\lesssim\|t|\xi_{1}|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t|\xi_{2}|\hat{v}\|_{L^{2}(\mathbb{R}^{2})}+\|t\partial_{t}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}$
and
$\|\widehat{\Omega_{12}v}\|_{L^{2}}\lesssim\left\|t|\xi_{2}|\partial_{t}\hat{v}\right\|_{L^{2}(\mathbb{R}^{2})}+\left\|t|\xi_{1}|\partial_{t}\hat{v}\right\|_{L^{2}(\mathbb{R}^{2})}\lesssim t^{1-\frac{2}{1+\varepsilon_{1}}}\tau\left\|v_{1}\right\|_{Z,1,\left(1+\varepsilon_{1},2\right)}.$
Collecting all the results above yields (4.3).
Next we estimate $\|\partial v\|_{Z,1,2}$. For this purpose, by an analogous analysis in the proof procedure of (3.52), it suffices only to treat $\|t|\xi|\partial_{t}\hat{v}\|_{L^{2}}$, $\||\xi_{j}||\xi|\hat{v}\|_{L^{2}}(j=1,2)$ and $\||\xi|(\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}})\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$.
It follows from (4.12)-(4.14) that
$\|t|\xi|\partial_{t}\hat{v}\|_{L^{2}}$ $\lesssim\big{\|}t^{1-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}|\xi|\hat{v}_{1}(\xi)\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}+\big{\|}t^{1-\frac{\mu}{2}}|\xi|^{\frac{\mu}{2}}\tau^{\mu}|\xi|\hat{v}_{1}(\xi)\big{\|}_{L^{2}(\tau|\xi|\leqslant 1,t|\xi|\geqslant 1)}+\left\|\tau|\xi|\hat{v}_{1}(\xi)\right\|_{L^{2}(t|\xi|\leqslant 1)}$ (4.21)
$\lesssim t^{-1}\tau\big{\|}t^{1-\frac{\mu}{2}}\tau^{\frac{\mu}{2}-1}t|\xi|\hat{v}_{1}\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}+t^{-1}\tau\big{\|}t^{1-\frac{\mu}{2}}\tau^{\frac{\mu}{2}-1}(\tau|\xi|)^{\frac{\mu}{2}}t|\xi|\hat{v}_{1}\big{\|}_{L^{2}(\tau|\xi|\leqslant 1,t|\xi|\geqslant 1)}$
$\quad+t^{-1}\left\|\tau\hat{v}_{1}(\xi)\right\|_{L^{2}(t|\xi|\leqslant 1)}$
$\lesssim t^{-1}\tau\|v_{1}\|_{Z,1,2}.$
By (4.6)-(4.8), one has that for $j=1,2$,
$\||\xi||\xi_{j}|\hat{v}\|_{L^{2}}$ $\lesssim\big{\|}t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}|\xi|^{-1}|\xi||\xi_{j}|\hat{v}_{1}\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}+\big{\|}t^{-\frac{\mu}{2}}|\xi|^{-\frac{\mu}{2}}\tau|\xi||\xi_{j}|\hat{v}_{1}\big{\|}_{L^{2}(\tau|\xi|\leqslant 1,t|\xi|\geqslant 1)}$ (4.22)
$\quad+\left\|\tau|\xi||\xi_{j}|\hat{v}_{1}\right\|_{L^{2}(t|\xi|\leqslant 1)}$
$\lesssim t^{-1}\tau\big{\|}t^{1-\frac{\mu}{2}}\tau^{\frac{\mu}{2}-1}|\xi_{j}|\hat{v}_{1}\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}+t^{-1}\tau\big{\|}(t|\xi|)^{1-\frac{\mu}{2}}|\xi_{j}|\hat{v}_{1}\big{\|}_{L^{2}(\tau|\xi|\leqslant 1,t|\xi|\geqslant 1)}$
$\quad+t^{-1}\tau\left\||\xi_{j}|\hat{v}_{1}(\xi)\right\|_{L^{2}(t|\xi|\leqslant 1)}$
$\lesssim t^{-1}\tau\|v_{1}\|_{Z,1,2}.$
As in (4.19), by (4.17), one obtains
$\||\xi|(\xi_{1}\partial_{\xi_{1}}\hat{v}+\xi_{2}\partial_{\xi_{2}})\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ (4.23)
$\lesssim\big{\|}t^{1-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}|\xi|\hat{v}_{1}(\xi)\big{\|}_{L^{2}(\tau|\xi|\geqslant 1)}+\|t^{-\frac{\mu}{2}}\tau^{\frac{\mu}{2}}\left(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}}\right)\hat{v}_{1}\left(\xi\right)\|_{L^{2}(\tau|\xi|\geqslant 1)}$
$\quad+\|t^{-\frac{\mu}{2}+1}\tau|\xi|^{-\frac{\mu}{2}+2}\hat{v}_{1}(\xi)\|_{L^{2}(\tau|\xi|\leq 1,t|\xi|\geq 1)}+t^{-\frac{\mu}{2}}\tau\|(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})\hat{v}_{1}(\xi)\|_{L^{2}(\tau|\xi|\leq 1,t|\xi|\geq 1)}$
$\quad+\|\tau|\xi|\hat{v}_{1}(\xi)\|_{L^{2}(t|\xi|\leq 1)}+t^{-1}\tau\|(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}})\hat{v}_{1}(\xi)\|_{L^{2}(t|\xi|\leq 1)}$
$\lesssim t^{-1}\tau\left\|t(|\xi_{1}|+|\xi_{2}|)\hat{v}_{1}\right\|_{L^{2}}+t^{-1}\tau\left\|\left(\xi_{1}\partial_{\xi_{1}}+\xi_{2}\partial_{\xi_{2}}\right)\hat{v}_{1}\right\|_{L^{2}}$
$\lesssim t^{-1}\tau\|v_{1}\|_{Z,1,2}.$
Therefore, in light of (4.21)-(4.23), then (4.4) and further Lemma 4.1 are proved. ∎
5 Proof of Theorems 1.1
Based on Lemmas 3.1-3.2 and Lemma 4.1, as in [21] or §$\S 4$ of [3], for any $T>1$, one can introduce the function space $X(T)$ with norm
$\|u\|_{X(T)}:=\sup_{t\in[1,T]}\big{(}t^{-(\delta-1)}\|u\|_{Z,1,2}+t\|\partial u\|_{Z,1,2}\big{)},$ (5.1)
where $\delta=\delta(\varepsilon_{1})$ is defined by Lemma 4.1.
Based on Duhamel’s principle and the expression (2.3), we define the nonlinear mapping $\mathcal{N}$ by
$\mathcal{N}u(t,x)$ $=\varepsilon\Psi_{0}(t,1,D)u_{0}(x)+\varepsilon\Psi_{1}(t,1,D)u_{1}(x)+\int_{1}^{t}\Psi_{1}(t,\tau,D)|u(\tau,x)|^{p}\mathrm{d}\tau$ (5.2)
$:=\varepsilon u^{lin}(t,x)+u^{non}(t,x)$
and introduce the following closed subset of $X(T)$
$X(T,M)=\left\{u\in X(T):\|u\|_{X(T)}\leq M\varepsilon\right\},$
where $M$ is a fixed positive constant to be determined (see (5.22) below). We now have
Proposition 5.1.
For any $\varepsilon_{1}\in(0,1)$, $T>1$ and $u,v\in X(T)$, the following estimates hold:
$\|\mathcal{N}u\|_{X(T)}\leq C\big{(}\varepsilon\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}+\|u\|_{X(T)}^{p}\big{)}$ (5.3)
and
$\|\mathcal{N}u-\mathcal{N}v\|_{X(T)}\leq C\|u-v\|_{X(T)}\big{(}\|u\|_{X(T)}^{p-1}+\|v\|_{X(T)}^{p-1}\big{)},$ (5.4)
where and below the constant $C>0$ is independent of $\varepsilon>0$ and $T>1$, and
$\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}=$ $\|u_{0}\|_{Z,1,2}+\|\nabla u_{0}\|_{Z,1,2}+\|u_{1}\|_{Z,1,2}+\|u_{1}\|_{Z,1,(1+\varepsilon_{1},2)}$ (5.5)
$+\|u_{0}+u_{1}\|_{Z,1,(1+\varepsilon_{1},2)}+\|u_{0}+u_{1}\|_{Z,1,2}.$
Proof.
By (3.2) and (4.3), one has
$\|\mathcal{N}u(t,\cdot)\|_{Z,1,2}$ $\lesssim t^{-\frac{\mu}{2}}\varepsilon\left\|u_{0}\right\|_{Z,1,2}+t^{\delta(\varepsilon_{1})-1}\varepsilon\left\|u_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta(\varepsilon_{1})-1}\varepsilon\left\|u_{0}+u_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}$ (5.6)
$\quad+t^{\delta(\varepsilon_{1})-1}\int_{1}^{t}\tau\left\||u(\tau,\cdot)|^{p}\right\|_{Z,1,(1+\varepsilon_{1},2)}\mathrm{d}\tau,$
where $\delta=\delta(\varepsilon_{1})$ is defined by Lemma 4.1. It follows from Lemma 2.1 in [21] that the following Hölder’s inequality holds
$\|fg\|_{(p,q)}\lesssim\|f\|_{(p_{1},q_{1})}\|g\|_{(p_{2},q_{2})},$ (5.7)
where $\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}\leq 1$ and $\frac{1}{q}=\frac{1}{q_{1}}+\frac{1}{q_{2}}\leq 1$. Then it follows from (5.7) and direct computation that
$\left\||u(\tau,\cdot)|^{p}\right\|_{Z,1,(1+\varepsilon_{1},2)}\lesssim\left\||u|^{p-1}\right\|_{(\tilde{q},2)}\|u(\tau,\cdot)\|_{Z,1,2}\leq\left\||u|^{p-1}\right\|_{(\tilde{q},\infty)}\|u(\tau,\cdot)\|_{Z,1,2},$ (5.8)
where $\tilde{q}=\tilde{q}(\varepsilon_{1})\in(2,\infty)$ satisfies
$\frac{1}{1+\varepsilon_{1}}=\frac{1}{2}+\frac{1}{\tilde{q}}.$ (5.9)
Denote $\kappa=\kappa(\varepsilon_{1})=\frac{2}{\tilde{q}}=\frac{1-\varepsilon_{1}}{1+\varepsilon_{1}}<1$. Then, by $p>2>1+\kappa$ and Hölder’s inequality (5.7), we can arrive at
$\left\||u(\tau,\cdot)|^{p-1}\right\|_{(\tilde{q},\infty)}\lesssim\|u(\tau,\cdot)\|_{\infty}^{p-1-\kappa}\|u(\tau,\cdot)\|_{(2,\infty)}^{\kappa}.$ (5.10)
In addition, the Sobolev embedding theorem on $S^{1}$ implies
$\|u(\tau,\cdot)\|_{(2,\infty)}^{2}\lesssim\int_{0}^{\infty}\left\|u(r\xi)\right\|_{H^{1}\left(S^{1}\right)}^{2}r\mathrm{d}r$ (5.11)
$\leq\int_{\mathbb{R}^{2}}u^{2}\mathrm{d}x_{1}\mathrm{d}x_{2}+\int_{0}^{\infty}\int_{0}^{2\pi}[\partial_{x_{1}}u(rcos\theta,rsin\theta)\cdot(-rsin\theta)+\partial_{x_{2}}u(rcos\theta,rsin\theta)\cdot(rcos\theta)]^{2}r\mathrm{d}r\mathrm{d}\theta$
$=\int_{\mathbb{R}^{2}}u^{2}\mathrm{d}x_{1}\mathrm{d}x_{2}+\int_{\mathbb{R}^{2}}(x_{1}\partial_{x_{2}}u-x_{2}\partial_{x_{1}}u)^{2}\mathrm{d}x_{1}\mathrm{d}x_{2}$
$\lesssim\|u\|_{Z,1,2}^{2}.$
On the other hand, it follows from the Klainerman-Sobolev inequality in Chapter 6 of [13] that
$\|u(\tau,\cdot)\|_{\infty}\lesssim\tau^{\frac{1}{2}}\left(\|u(\tau,\cdot)\|_{Z,1,2}+\|\partial u(\tau,\cdot)\|_{Z,1,2}\right).$ (5.12)
Substituting (5.11) and (5.12) into (5.10) yields
$\left\||u(\tau,\cdot)|^{p-1}\right\|_{(\tilde{q},\infty)}\lesssim\tau^{\frac{p-1-\kappa}{2}}\left(\|u(\tau,\cdot)\|_{Z,1,2}+\|\partial u(\tau,\cdot)\|_{Z,1,2}\right)^{p-1-\kappa}\|u\|_{Z,1,2}^{\kappa}.$
By the definition of norm (5.1) in space $X(T)$, it holds that
$\left\||u(\tau,\cdot)|^{p-1}\right\|_{(\tilde{q},\infty)}\lesssim\tau^{\frac{p-1-\kappa}{2}+(p-1-\kappa)(\delta-1)+\kappa(\delta-1)}\|u\|_{X(T)}^{p-1}.$
Substituting it into (5.8) and then (5.6) implies
$\|\mathcal{N}u(t,\cdot)\|_{Z,1,2}$ $\lesssim t^{-\frac{\mu}{2}}\varepsilon\left\|u_{0}\right\|_{Z,1,2}+t^{\delta-1}\varepsilon\left\|u_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+t^{\delta-1}\varepsilon\left\|u_{0}+u_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}$ (5.13)
$\quad+t^{\delta-1}\|u\|_{X(T)}^{p}\int_{1}^{t}\tau^{1+\frac{p-1-\kappa}{2}+(p-1-\kappa)(\delta-1)+(\kappa+1)(\delta-1)}\mathrm{d}\tau,$
Note that $\delta=\frac{2\varepsilon_{1}}{1+\varepsilon_{1}}\rightarrow 0+$ and $\kappa=\frac{1-\varepsilon_{1}}{1+\varepsilon_{1}}\rightarrow 1-$ as $\varepsilon_{1}\rightarrow 0+$. Therefore, for $p>2$, one can choose a sufficiently small $\varepsilon_{1}>0$ such that
$11+\frac{p-1-\kappa}{2}+p(\delta-1)<-$ 1 .
Then
$t^{-(\delta-1)}\|\mathcal{N}u(t,\cdot)\|_{Z,1,2}$ $\lesssim\varepsilon\left\|u_{0}\right\|_{Z,1,2}+\varepsilon\left\|u_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+\varepsilon\left\|u_{0}+u_{1}\right\|_{Z,1,(1+\varepsilon_{1},2)}+\|u\|_{X(T)}^{p}$ (5.14)
$\lesssim\varepsilon\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}+\|u\|_{X(T)}^{p}.$
Similarly, by (3.52) and (4.4), we have
$\|\partial\mathcal{N}u(t,\cdot)\|_{Z,1,2}$ $\lesssim t^{-\frac{\mu}{2}}\varepsilon\left\|\nabla u_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\varepsilon\left\|u_{1}\right\|_{Z,1,2}+t^{-1}\varepsilon\left\|u_{0}+u_{1}\right\|_{Z,1,2}$ (5.15)
$\quad+t^{-1}\int_{1}^{t}\tau\left\||u(\tau,\cdot)|^{p}\right\|_{Z,1,2}\mathrm{d}\tau.$
As treated in (5.8), one has
$\left\||u(\tau,\cdot)|^{p}\right\|_{Z,1,2}\lesssim\left\||u(\tau,\cdot)|^{p-1}\right\|_{2+\varepsilon_{2}}\|u(\tau,\cdot)\|_{Z,1,\bar{q}},$ (5.16)
where $\bar{q}=\bar{q}(\varepsilon_{2})\in(2,\infty)$ fulfills
$\frac{1}{2}=\frac{1}{2+\varepsilon_{2}}+\frac{1}{\bar{q}}.$ (5.17)
In addition, it follows from Sobolev embedding theorem that
$\|u(\tau,\cdot)\|_{Z,1,\bar{q}}\lesssim\|u(\tau,\cdot)\|_{Z,1,2}+\|\partial u(\tau,\cdot)\|_{Z,1,2}.$ (5.18)
Moreover, according to $p>2$, one has
$\left\||u(\tau,\cdot)|^{p-1}\right\|_{2+\varepsilon_{2}}\leq\|u(\tau,\cdot)\|_{\left(2+\varepsilon_{2}\right)(p-1)}^{p-1}$ $\lesssim\tau^{-\left(\frac{1}{2}-\frac{1}{\left(2+\varepsilon_{2}\right)(p-1)}\right)(p-1)}\|u(\tau,\cdot)\|_{Z,1,2}^{p-1}$
$\leq\|u(\tau,\cdot)\|_{Z,1,2}^{p-1}.$
This, together with (5.18) and (5.16), yields
$\|\partial\mathcal{N}u(t,\cdot)\|_{Z,1,2}$ $\lesssim t^{-\frac{\mu}{2}}\varepsilon\left\|\nabla u_{0}\right\|_{Z,1,2}+t^{-\frac{\mu}{2}}\varepsilon\left\|u_{1}\right\|_{Z,1,2}+t^{-1}\varepsilon\left\|u_{0}+u_{1}\right\|_{Z,1,2}$
$\quad+t^{-1}\|u\|_{X(T)}^{p}\int_{1}^{t}\tau^{1+p(\delta-1)}\mathrm{d}\tau.$
For $p>2$, we choose a fixed small constant $\varepsilon_{1}>0$ such that $11+p(\delta(\varepsilon_{1})-1)<-$ 1 holds. Thus,
$t\|\partial\mathcal{N}u(t,\cdot)\|_{Z,1,2}$ $\lesssim\varepsilon\left\|\nabla u_{0}\right\|_{Z,1,2}+\varepsilon\left\|u_{1}\right\|_{Z,1,2}+\varepsilon\left\|u_{0}+u_{1}\right\|_{Z,1,2}+\|u\|_{X(T)}^{p}$ (5.19)
$\lesssim\varepsilon\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}+\|u\|_{X(T)}^{p}.$
Combining (5.14) and (5.19) derives (5.3).
We next prove the contractive estimate (5.4). As treated in (5.6), we have that for $u,v\in X(T)$ and $p>2$,
$\|(\mathcal{N}u-\mathcal{N}v)(t,\cdot)\|_{Z,1,2}$ $\lesssim t^{\delta-1}\int_{1}^{t}\tau\left\||u-v|(\tau,\cdot)(|u|+|v|)^{p-1}(\tau,\cdot)\right\|_{Z,1,(1+\varepsilon_{1},2)}\mathrm{d}\tau$ (5.20)
$\lesssim t^{\delta-1}\int_{1}^{t}\tau\left\|(|u|+|v|)^{p-1}(\tau,\cdot)\right\|_{(\tilde{q},\infty)}\||u-v|(\tau,\cdot)\|_{Z,1,2}\mathrm{d}\tau$
$\lesssim t^{\delta-1}\|u-v\|_{X(T)}\big{(}\|u\|_{X(T)}^{p-1}+\|v\|_{X(T)}^{p-1}\big{)}\int_{1}^{t}\tau^{1+\frac{p-1-\kappa}{2}+p(\delta-1)}\mathrm{d}\tau$
$\lesssim t^{\delta-1}\|u-v\|_{X(T)}\big{(}\|u\|_{X(T)}^{p-1}+\|v\|_{X(T)}^{p-1}\big{)},$
where $\tilde{q}$ is defined by (5.9). On the other hand, employing the same approach for deriving (5.19), one can obtain that for $u,v\in X(T)$ and $p>2$,
$\|\partial(\mathcal{N}u-\mathcal{N}v)(t,\cdot)\|_{Z,1,2}$ $\lesssim t^{-1}\int_{1}^{t}\tau\left\||u-v|(\tau,\cdot)(|u|+|v|)^{p-1}(\tau,\cdot)\right\|_{Z,1,2}\mathrm{d}\tau$ (5.21)
$\lesssim t^{-1}\int_{1}^{t}\tau\left\|(|u|+|v|)^{p-1}(\tau,\cdot)\right\|_{{}_{2+\varepsilon_{1}}}\||u-v|(\tau,\cdot)\|_{Z,1,\bar{q}}\mathrm{d}\tau$
$\lesssim t^{-1}\|u-v\|_{X(T)}\big{(}\|u\|_{X(T)}^{p-1}+\|v\|_{X(T)}^{p-1}\big{)}\int_{1}^{t}\tau^{1+p(\delta-1)}\mathrm{d}\tau$
$\lesssim t^{-1}\|u-v\|_{X(T)}\big{(}\|u\|_{X(T)}^{p-1}+\|v\|_{X(T)}^{p-1}\big{)},$
where $\bar{q}$ is defined by (5.17). Therefore, (5.4) is shown and then Proposition 5.1 is established by collecting (5.1), (5.20)-(5.21) and (5.3).
Based on Proposition 5.1, we start to prove Theorem 1.1.
Proof of Theorem 1.1.
Choosing $M=3C\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}$ in $X(T,M)$, where $C$ is the positive constant mentioned in Proposition 5.1. Then for any $u\in X(T,M)$, one can obtain
$\|\mathcal{N}u\|_{X(T)}$ $\leq C\varepsilon\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}+C\|u\|_{X(T)}^{p}$ (5.22)
$\leq C\varepsilon\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}+C\left(3C\varepsilon\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}\right)^{p}$
$\leq\left(C+3^{p}C^{p+1}\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}^{p-1}\varepsilon^{p-1}\right)\varepsilon\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}$
$\leq 3C\varepsilon\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}},$
where $\varepsilon\leq\big{(}\frac{2}{3^{p}C^{p}}\big{)}^{\frac{1}{p-1}}\frac{1}{1+\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}}$. Similarly,
$\|\mathcal{N}u-\mathcal{N}v\|_{X(T)}$ $\leq C\|u-v\|_{X(T)}\left(\|u\|_{X(T)}^{p-1}+\|v\|_{X(T)}^{p-1}\right)$ (5.23)
$\leq C\|u-v\|_{X(T)}\cdot 2\left(3C\varepsilon\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}\right)^{p-1}$
$\leq\frac{1}{2}\|u-v\|_{X(T)},$
provided that $\varepsilon\leq\frac{{(4C)}^{-\frac{1}{p-1}}}{3C(1+\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}})}$. Therefore, by setting $\varepsilon_{0}:=\min\{\big{(}\frac{2}{3^{p}C^{p}}\big{)}^{\frac{1}{p-1}}\frac{1}{1+\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}}},$ $\frac{{(4C)}^{-\frac{1}{p-1}}}{3C(1+\left\|\left(u_{0},u_{1}\right)\right\|_{\mathcal{A}})}\}$ and $\varepsilon\leq\varepsilon_{0}$, one can conclude that $\mathcal{N}$ is a contractive mapping from $X(T,M)$ into $X(T,M)$. Therefore, there exists a unique solution $u\in X(T,M)$ such that $\mathcal{N}u=u$. This yields that $u$ is a global solution of (1.1) since the related constant $C>0$ is independent of $T$. Then the proof of Theorem 1.1 is completed. ∎
Acknowledgements. Yin Huicheng wishes to express his deep gratitude to Professor Ingo Witt (University of Göttingen, Germany) for his constant interests in this problem and many fruitful discussions in the past. In addition, the authors would like to thank Dr. He Daoyin for his invaluable discussions.
Conflict of Interest Statement:
The authors declare that there is no conflict of interest in relation to this article.
Data availability statement:
Data sharing is not applicable to this article as no data sets are generated during the current study.

### v2

@@
 In our previous papers, the global small solution $u$ has been obtained for $p>p_{s}(2+\mu)$ and $20<\mu<$ 2 but $\mu\not=1$.
-In the present paper, by the vector field method together with the delicate analysis on the Bessel functions, we will show the global existence of small solution $u$ for $p>2$ and $\mu\geq 2$.
+In the present paper, by the vector field method together with the delicate analysis on the Bessel functions, we will show the global existence of small solution $u$ for $p>2$ and $\mu>2$.
 In forthcoming paper, for $\mu=1$ and $p>p_{s}(2+\mu)=p_{s}(3)=1+\sqrt{2}$, the global solution $u$ is also obtained.
-Therefore, this open question has been solved completely.
+Therefore, collecting our series of conclusions together with partial results from others, this open question has been solved completely.
 Keywords. Scale-invariant damping, global existence, Bessel function, vector field,
@@
 In our forthcoming paper [12], (A1) with $\mu=1$ and $p>p_{s}(2+\mu)=1+\sqrt{2}$ will be shown by applying the Bessel function tools and deriving some suitable time-decay or spacetime-decay estimates of weak solution $u$ as in [1], [11] and [19].
-Collecting the results in [19] and [11] (for $20<\mu<$ 2 but $\mu\not=1$, $p>p_{s}(2+\mu)$), [12] (for $\mu=1$, $p>p_{s}(2+\mu)=1+\sqrt{2}$) together with Theorem 1.1 in the paper (for $32<\mu<$ 3 and $p\geq 2$) , [3, Theorem 2] (for $\mu=2$ and $p\geq 2$) and [2, Theorem 2] (for $\mu\geq 3$ and $p\geq 2$), then the Open question (A) has been solved completely.
+Collecting the results in [19] and [11] (for $20<\mu<$ 2 but $\mu\not=1$, $p>p_{s}(2+\mu)$), [12] (for $\mu=1$, $p>p_{s}(2+\mu)=1+\sqrt{2}$) together with Theorem 1.1 in the paper (for $32<\mu<$ 3 and $p>2$) , [3, Theorem 2] (for $\mu=2$ and $p>2$) and [2, Theorem 2] (for $\mu\geq 3$ and $p>2$), then the Open question (A) has been solved completely.
 Remark 1.2.
@@
 $\square Z^{I}\phi+\sum_{i,j,k=0}^{n}\sum_{|I^{\prime}|+|I^{\prime\prime}|\leq|I|}C_{ij}^{k}\partial_{ij}^{2}Z^{I^{\prime}}\phi\partial_{k}Z^{I^{\prime\prime}}\phi=0.$ (1.3)
-Therefore, it follows from (1.3) that $\{Z^{I}\phi\}_{|I|\leq N_{0}}$ consist of a closed nonlinear wave system and then $\|Z^{I}\partial\phi\|_{L_{x}^{2}(\mathbb{R}^{n})}$ can be obtained by the standard energy estimate.
+Therefore, it follows from (1.3) that $\{Z^{I}\phi\}_{|I|\leq N_{0}}$ consists of a closed nonlinear wave system and then $\|Z^{I}\partial\phi\|_{L_{x}^{2}(\mathbb{R}^{n})}$ can be obtained by the standard energy estimate.
 By the crucial Klainerman-Sobolev inequality
@@
 However, for the general 2-D equation $\square u+\frac{\mu}{t}\,\partial_{t}u=|u|^{p}$ with $32<\mu<$ 3, it is difficult to find a suitable transformation of the unknown function $u$ such that a related semilinear wave equation is derived.
-Moreover, the commutator $[\square+\frac{\mu}{t}\,\partial_{t},Z]$ can not be expressed an efficient linear combination of $\square+\frac{\mu}{t}\,\partial_{t}$ and $Z$.
+Moreover, the commutator $[\square+\frac{\mu}{t}\,\partial_{t},Z]$ can not be expressed as an efficient linear combination of $\square+\frac{\mu}{t}\,\partial_{t}$ and $Z$.
 Indeed, letting $w_{1}=\partial_{t}u$ and $w_{2}=L_{1}u$, then one has from $\square u+\frac{\mu}{t}\,\partial_{t}u=|u|^{p}$ that
@@
 and
-$\square w_{2}+\frac{\mu}{t}\,\partial_{t}w_{2}-\frac{\mu x_{1}}{t^{2}}w_{1}=L_{1}(|u|^{p}).$ (1.6)
-This means that the linear parts in the nonlinear equations (1.5) and (1.6) become more complicated Euler-Poisson-Darboux-Klein-Gordon’s type models.
+$\square w_{2}+\frac{\mu}{t}\,\partial_{t}w_{2}-\frac{\mu}{t}\partial_{1}u-\frac{\mu x_{1}}{t^{2}}w_{1}=L_{1}(|u|^{p}).$ (1.6)
+This means that the linear parts in the nonlinear equations (1.5) and (1.6) become more complicated than the Euler-Poisson-Darboux operator together with the Klein-Gordon type’s operator.
 To overcome this essential difficulty, we will take the following measures.
@@
 For $|\xi|\leq 1$ and $t|\xi|\geq 1$,
-$\left|\partial_{t}^{2}\Psi_{0}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{3}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho-1}=(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|,$ (3.59)
-$\left|\partial_{t}\Psi_{1}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{2}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho}=(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|.$
+$\left|\partial_{t}^{2}\Psi_{0}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{3}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho-1}\leq(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|,$ (3.59)
+$\left|\partial_{t}^{2}\Psi_{1}(t,1,\xi)\right|\lesssim t^{\rho}|\xi|^{2}(t|\xi|)^{-\frac{1}{2}}|\xi|^{\rho}\leq(t|\xi|)^{-\frac{\mu}{2}}\cdot|\xi|.$
 For $t|\xi|\leq 1$,
 $\left|\partial_{t}^{2}\Psi_{0}(t,1,\xi)\right|\lesssim t^{-1}|\xi|,$ (3.60)
-$\left|\partial_{t}\Psi_{1}(t,1,\xi)\right|\lesssim t^{-1}|\xi|.$
+$\left|\partial_{t}^{2}\Psi_{1}(t,1,\xi)\right|\lesssim t^{-1}|\xi|.$
 By comparing the estimates in (3.58), (3.59) and (3.60) with the corresponding ones in (3.16), (3.20) and (3.24), it is known that the treatment of $\|\partial_{t}^{2}\hat{v}\|_{L^{2}(\mathbb{R}^{2})}$ can be similar to the one for $\left\||\xi|\partial_{t}\hat{v}(t,\cdot)\right\|_{L^{2}(\mathbb{R}^{2})}$.
@@
 $\left|\partial_{t}\Psi_{1}(t,\tau,\xi)\right|\lesssim|\xi|t^{\rho}\tau^{1-\rho}(\tau|\xi|)^{-\rho}(t|\xi|)^{-\frac{1}{2}}=t^{-\frac{\mu}{2}}|\xi|^{\frac{\mu}{2}}\tau^{\mu}.$ (4.13)
-For $\tau|\xi|\leq t|\xi|\leq 1$, by (2.8) and (2.25) it holds
+For $\tau|\xi|\leq t|\xi|\leq 1$, by (2.8) and (2.25), it holds
 $\left|\partial_{t}\Psi_{1}(t,\tau,\xi)\right|$ $\lesssim|\xi|t^{\rho}\tau^{1-\rho}\left[(\tau|\xi|)^{-\rho}(t|\xi|)^{\rho-1}+(\tau|\xi|)^{\rho}(t|\xi|)^{1-\rho}\right]$ (4.14)
@@
 Therefore, in light of (4.21)-(4.23), then (4.4) and further Lemma 4.1 are proved. ∎
-5 Proof of Theorems 1.1
+5 Proof of Theorem 1.1
 Based on Lemmas 3.1-3.2 and Lemma 4.1, as in [21] or §$\S 4$ of [3], for any $T>1$, one can introduce the function space $X(T)$ with norm
@@
 $\lesssim\|u\|_{Z,1,2}^{2}.$
-On the other hand, it follows from the Klainerman-Sobolev inequality in Chapter 6 of [13] that
+On the other hand, it follows from the Klainerman-Sobolev inequality in [16] that
 $\|u(\tau,\cdot)\|_{\infty}\lesssim\tau^{\frac{1}{2}}\left(\|u(\tau,\cdot)\|_{Z,1,2}+\|\partial u(\tau,\cdot)\|_{Z,1,2}\right).$ (5.12)